### v1

# Retain or Reframe? A Computational Framework for the Analysis of Framing in News Articles and Reader Comments

Abstract
When a news article describes immigration as an "economic burden" or a "humanitarian crisis," it selectively emphasizes certain aspects of the issue. Although framing shapes how the public interprets such issues, audiences do not absorb frames passively but actively reorganize the presented information. While this relationship between source content and audience response is well-documented in the social sciences, NLP approaches often ignore it, detecting frames in articles and responses in isolation. We present the first computational framework for large-scale analysis of framing across source content (news articles) and audience responses (reader comments). Methodologically, we refine frame labels and develop a framework that reconstructs dominant frames in articles and comments from sentence-level predictions, and aligns articles with topically relevant comments. Applying our framework across eleven topics and two news outlets, we find that frame reuse in comments correlates highly across outlets, while topic-specific patterns vary. We release a frame classifier that performs well on both articles and comments, a dataset of article and comment sentences manually labeled for frames, and a large-scale dataset of articles and comments with predicted frame labels. [^1] We released our at: https://anonymous.4open.science/r/FrAC-A8E3/; and fine-tuned frame classifier: https://huggingface.co/mattdr/sentence-frame-classifier
1 Introduction
Framing research has long recognized that how news content is presented fundamentally shapes public understanding of issues (Entman, 1993; Chong and Druckman, 2007). However, audiences do not passively absorb frames but rather "actively filter, sort and reorganize information in personally meaningful ways in the process of constructing an understanding of public issues" (Neuman et al., 1992, pg.77). Such a constructionist understanding emphasizes that frames cannot be fully understood without considering the interaction between source and receiver (Scheufele, 1999; Gorp, 2007; Hubner and Dixon, 2023). This is particularly evident on news websites with comment threads that create immediate opportunities for readers to publicly respond to framed content — a paradigm shift from one-way communication to participatory discourse where audiences actively engage with and reconstruct news content (Risch and Krestel, 2020).
However, current research has yet to fully explore the relationship between media framing and audience reception. Cognitive studies have demonstrated that the framing of news articles can significantly shape readers’ perceptions of an issue in controlled, small-scale experiments (Price et al., 1997; Valkenburg et al., 1999; Nelson and Oxley, 1999). Communication scholars have examined the influence of framing by analyzing online news articles and their associated reader comments (Zhou and Moy, 2007; Muñiz et al., 2015; Dargay, 2016; Hubner and Dixon, 2023). These studies often focused on a single issue and relied on small, manually annotated data sets, which limited generalizability and led to inconsistent findings.
Meanwhile, Natural Language Processing (NLP) approaches have developed numerous classification systems (for a review, see Ali and Hassan (2022)) aimed at detecting frames in text of different lengths such as news articles (Field et al., 2018), or social media posts (Mendelsohn et al., 2021). Yet these large-scale methods treat frame detection in source documents (such as news) and responses to them (such as comments or social media posts) as isolated classification tasks without connecting the framed source text and the framing in the audience response (Otmakhova et al., 2024).
Figure 1: Overview of our framework. Bottom: Article with two comments illustrating reframing (Comment 1; changes the article frame) and frame retention (Comment 2; keeps the article frame).
To address this gap, we propose a computational framework that investigates at scale the relationship between how information is framed in news articles and how readers respond to such framed content in associated comments. Drawing from the constructionist understanding of framing (Gorp, 2007) and using an established framing taxonomy (Boydstun et al., 2014; Card et al., 2015), we examine both source content and audience responses as interconnected elements of a broader communicative process (Hubner and Dixon, 2023).
Throughout this paper, we use media framing effects to describe the relation between news article framing and that of reader comments, as an umbrella term that covers two phenomena: frame retention occurs when readers adopt the dominant frame used in the source article in their comments, while reframing occurs when readers shift to a different dominant frame in their comments. Figure 1 illustrates both, with an example.
Our large-scale analysis across eleven diverse topics and two news outlets [^2] The New York Times (USA) and The Globe and Mail (Canada) reveals systematic patterns of media framing effects. We find that (a) frame retention rates vary by source document frame, and this variation correlates highly across the two outlets, with value-laden frames being more often reframed; (b) retention patterns vary systematically by topic, and (c) when reframing, readers predominantly shift discussions toward political interpretation regardless of the original article frame.
We make three main contributions: (1) a computational framework for investigating the relationship between source frames (articles) and frames in associated responses (comments), (2) a frame classification model that effectively generalizes across different text types (news articles and reader comments) and topics, and (3) two data sets: a manually labeled evaluation corpus (FrAC) and a large-scale data set of articles and comments aligned by topic group and labeled with dominant frames, for reproducibility and supporting future research.
2 Related Work
Cognitive Studies on Framing Effects
Cognitive research established that news frames significantly influence audience interpretation using controlled experimental designs, where participants were assigned to different framing conditions of the same issue and asked to report their thoughts. Valkenburg et al. (1999) examined how news stories framed as conflict, human interest, attribution of responsibility, or economic consequences affected readers, finding that participants generally adopted news frames. However, Price et al. (1997) showed that frames not only guide interpretation but also stimulate new thoughts and feelings not explicitly present in the original content. Nelson and Oxley (1999) and Shen (2004) showed that framing primarily influences the importance of existing beliefs rather than changing the beliefs themselves, with frames most effective when resonating with pre-existing schemas. While producing valuable insights, these studies were restricted to small-scale experimental settings and narrow demographic representation. This limited their generalizability to framing effects in natural media environments with direct audience responses.
Media Framing Effects in Communication Studies
Communication researchers have attempted to bridge this gap by analyzing reader comments on online news articles, albeit with mixed results. Several studies have found evidence of frame retention. Kleut and Milojevic (2021) analyzed online coverage of Serbian protests and showed that comments were influenced by the article’s frames. Dargay (2016) examined terrorism coverage in The New York Times, finding that commentators engaged with dominant frames of the articles, although they also introduced new frames.
Other studies have found no evidence of retention. Holton et al. (2014) analyzed online health coverage across major outlets and found no evidence of frame retention in comment threads, noting a disconnect between thematic news frames and predominantly episodic comment frames. Muñiz et al. (2015) examined coverage of an oil company expropriation in Spanish online outlets and found that media frames did not produce corresponding effects in reader comments. Similarly, Zhou and Moy (2007) found no framing effects in news about Chinese sociopolitical incidents. Finally, in climate change coverage, Koteyko et al. (2013) found minimal retention of frames in reader comments on online articles from UK outlets, while Hubner and Dixon (2023) showed that retweets sharing climate news rarely replicated the original frames.
There are several explanations for these inconsistencies, including small data sets, focus on a single topic, and the different adopted framing conceptualizations. In this study, we increase the data set size, keep the framing conceptualization fixed but apply it across eleven topics, in order to disentangle some of these confounds.
Computational Approaches
NLP approaches to framing have evolved from exploratory topic-modeling (DiMaggio et al., 2013) to structured classification tasks (for a review, see Ali and Hassan (2022)). The Media Frame Corpus (MFC, Card et al. (2015)), based on Boydstun et al. (2014), has been particularly influential, spurring various frame classification methods including probabilistic approaches (e.g. Johnson and Goldwasser (2018)), neural networks (e.g. Naderi and Hirst (2017)), and pre-trained language models (e.g. Khanehzar et al. (2021)). These computational approaches have analyzed frames at different granularities: headline level (Liu et al., 2019a; Akyürek et al., 2020), document level (Field et al., 2018; Ji and Smith, 2017; Piskorski et al., 2023), span level (Khanehzar et al., 2021), sentence level (Naderi and Hirst, 2017; Hartmann et al., 2019), and paragraph level (Roy and Goldwasser, 2020). Some studies have investigated the generalizability of the MFC to other domains and languages: Khanehzar et al. (2019) found reduced performance when applying fine-tuned BERT-based classifiers to different English contexts (Australian parliamentary speeches), while Daffara et al. (2025) tested cross-linguistic transfer using both fine-tuned pre-trained language models and zero-shot LLM classification on Portuguese news articles. However, these studies either focused on cross-linguistic transfer, or on the diagnosis of the problem of cross-topic and genre generalization but fell short of offering a solution to the problem, which we do here.
3 Research Questions
The gaps in existing research motivate three fundamental questions about media framing effects in readers comments:
  1. 1.
    To what extent are frames retained vs changed? Previous studies have shown mixed results. We investigate at scale whether article frames consistently influence comment frames, and whether this influence varies systematically across different types of frames.
  2. 2.
    How do framing effects vary across topics? Topic-specific studies cannot reveal whether framing effects represent universal phenomena or topic-specific patterns. We examine whether framing effects are consistent or vary based on the topic being discussed.
  3. 3.
    When readers do not retain the dominant frame of the article, how do they reframe it? Understanding reframing patterns is crucial for comprehending how audiences actively reconstruct media messages. We examine systematic tendencies in how readers shift framing when they reject source frames.
4 Framework
To address these research questions, we propose a scalable and automated framework for analyzing media framing effects (retention and reframing).
4.1 Data
Our framework requires a data set of source documents $d_{i}$ (e.g. news articles, headlines, tweets, etc.) each paired with a set of response documents $R_{d_{i}}=\{r_{1},r_{2},\ldots\}$ (e.g. comments, retweets, blog posts, etc.) directly linked to the source ($D=\{(d_{i},R_{d_{i}})\}_{i=1}^{N}$). For the purposes of this study, we use two existing data sets of news articles with their associated comments:
  • SOCC (The SFU Opinion Comments Corpus; Kolhatkar et al. (2020)) [^3] Available at: https://www.kaggle.com/datasets/mtaboada/sfu-opinion-and-comments-corpus-socc: consists of 10,339 articles and over 1.2 million comments sourced from The Globe and Mail, a major Canadian news paper. The corpus covers five years (2012–2016).
  • NYT (New York Times): consists of articles and comments from The New York Times [^4] Available at: https://www.kaggle.com/datasets/aashita/nyt-comments. Article texts were retrieved using the article titles in the data set via ProQuest. with over two million comments associated with approximately 9,000 articles published between January-May 2017 and January-April 2018.
We note that the framework generalizes to any corpus where source documents are mapped to audience responses.
4.2 Framework Structure
The framework consists of three core components that work in concert to analyze framing dynamics between articles and comments (Figure 1):
  1. 1.
    Frame Classification: To compare framing across source documents and responses, we build on established framing taxonomies to develop and evaluate frame classifiers that work reliably across both document types.
  2. 2.
    Dominant Frame Detection: In line with prior work in cognitive and communication studies (Section 2), we identify and compare a single, dominant frame (defined as the most recurring one; Section 6.2) in source documents and responses. We thus require a method to determine the dominant frame in each input document.
  3. 3.
    Topic Identification: We assign each source document and reader response to a topic, for two reasons. First, assigning topics to articles allows us to perform analysis of framing effects over multiple topics, advancing over previous, topic-specific studies (Section 2) and potentially explaining their divergent results. Second, we want to capture framing effects only for responses where commentators stay on topic with the source, rather than digress into a different topic (with potentially a different frame, which can lead to underestimation of retention rates).
Below, we first introduce a new corpus that we annotated for the frame classification subtask (Section 5). Section 6 explains and evaluates our implementations of the three framework components introduced above. Finally, we apply the framework to construct a data set (Section 6.4) that we use to analyze media framing effects (Section 7).
5 The FrAC Corpus
To apply our framework, we require (1) a frame label set that works robustly across both source documents (news articles) and reader reactions (comments), and across a variety of different topics; and (2) a way to validate predicted frames across these two types of documents. To this end, we introduce the Frames in Articles and Comments (FrAC) corpus. Our corpus uses a more robust frame label set derived from that of the Media Frames Corpus (Section 5.1), and has high-quality manual annotations showing that humans can reliably apply our label set to both document types (Section 5.2).
5.1 Frame Labels and Granularity
We devise a robust frame label set, building on the Media Frames Corpus (MFC). The MFC is a foundational data set in computational framing analysis based on Boydstun et al. (2014) theoretical framework operationalizing framing through 15 topic-agnostic frame categories. The original MFC corpus Card et al. (2015) consists only of news articles; however, more recently the framework was successfully applied to other types of documents, including online forum comments Hartmann et al. (2019), suggesting suitability for our task that involves both long, formal documents and short, less formal responses.
However, it has been noted (e.g., Ali and Hassan (2022); Daffara et al. (2025)) that some MFC frames are poorly defined and difficult to distinguish, leading to low inter-annotator agreement and model performance. First, to apply the MFC labels across our two text types (articles and comments), we need to define a granularity that is meaningful across both of them. While most of NLP studies on MFC looked at article-level dominant frames in the MFC, it is not feasible for comments (or other response documents), which are often brief and composed of single sentences. We therefore chose to use the sentence level – the smallest shared textual unit between the two types of text. [^5] Section 6.2 explains how we derive a dominant document frame from sentence-level frames.
Second, to improve robustness, we isolated highly confused frame labels in the original MFC data set. To do so, we first identified spans-level MFC annotations (of ${\geq}3$ words) to which two or more annotators assigned different frame labels. We examined such inconsistent annotations using confusion matrices – both for all MFC data and separately for each topic included in the MFC corpus. We took a conservative approach by only merging frame labels that were confused consistently across all topics, [^6] For example, “Health and Safety” and “Legality, constitutionality and jurisprudence” labels are frequently assigned to the same spans in “Tobacco” and “Gun Control”, but not in other topics. In this case, we conclude that the confusion is due to the topic specifics rather than meaning of the label, and keep such semantically distinct labels separate. resulting in two merges of two frames (Table 1). To further improve robustness of the label set, we removed the two least frequent classes (“External regulation and reputation” and “Capacity and resources”) which accounted for less than 0.7% of the data. We finally obtained 9 frame categories, plus a generic “Other” frame to account for instances that do not match any of our nine frames. Table 1 lists the frames, with full definitions in Appendix Table 8.
1) Legality and Crime, 2) Political and Policies+, 3) Economic, 4) Health and Safety, 5) Cultural Identity, 6) Public Opinion, 7) Morality, 8) Fairness and Equality, 9) Security and Defense, 10) Other.
Table 1: The FrAC frame label set. Merged MFC labels "Legality, constitutionality and jurisprudence" and "Crime and punishment". + Merged MFC labels "Policy prescription and evaluation" and "Political".
5.2 Human Annotation
To validate our frame label set (and model predictions) we collect human annotation on a subset of the SOCC corpus (Section 4.1).
Annotation Setup
A custom web-based annotation interface was developed using FastAPI. [^7] https://annotation-app-slty.onrender.com/. The interface included detailed instructions and examples, and an interactive training phase with immediate feedback and explanations. See Appendix 5 for detailed descriptions.
We recruited six native English speakers as annotators, all with a background in NLP or political science and journalism. [^8] This study was approved by Human Ethics Committee (Referance No. 2025-32561-67793-4), and has been taken out to the according ethical standards. Annotators were compensated with vouchers over 50 USD, well above local minimum wage requirements. Each annotator was presented with 100 sentences from the SOCC corpus, drawn evenly from articles and comments, 20% of the sentences overlapped with another annotator to assess agreement. One additional annotator (an author of this paper) annotated all sentences (so all sentences in the resulting data set had at least two annotations). We adopted a multi-label annotation setup. Annotators were presented one input sentence at a time, and were asked to select at least one out of our 9 frames. If no frame matched, the category “Other” should be selected.
Inter-Annotator Agreement
As annotators could assign multiple labels to each sentence, we assessed inter-annotator agreement in two ways. Following prior work Arora et al. (2025), we use label-wise Krippendorff’s $\alpha$ to assess per-label agreement, and the Jaccard index to measure the averaged overlap in label sets between annotators. The averaged Krippendorff’s $\alpha$ across all labels and annotators indicates moderate to high agreement ($\alpha=0.76$), with some variance across frame labels (min for Morality with $\alpha=0.63$, max for Economic with $\alpha=0.89$). However even the category with lowest $\alpha$ reaches moderate agreement. The Jaccard index of 0.79 confirms this trend, suggesting a substantial overlap across label sets. This shows that a more robust set of labels applied at a sentence level allows to achieve substantial agreement despite the purported subjectivity of the task.
Gold Standard
To construct a gold standard for model evaluation, we applied majority voting among annotators for each sentence. A clear consensus (where either two or three annotators chose the same label) was achieved for 532 cases (where in 500 cases a single label was assigned, while in 32 all annotators have chosen two consistent labels). Sentences where annotators disagreed (45) were adjudicated by two authors of the paper, who independently re-assessed the cases of contention, choosing the label out of 9 categories. We kept only the cases where both adjudicators agreed with one of the original annotations, discarding 13 sentences. This process resulted in a validation data set of 556 sentences, which we use in Section 6.1 to assess model performance.
6 Methodology
Having established a general framework, and a conceptualization of media frames that generalize across articles and comments, we next describe the implementation of the framework components on our SOCC and NYT data sets of articles and comments (Section 4.1).
6.1 Frame Classification
We train frame classifiers that predict sentence-level media frames for both comments and articles, following the framing schema defined in Section 5.1. To do so, we require suitable training data from both genres, and a robust evaluation setup.
Frame Classification Data
We source articles from the Media Frames Corpus (Card et al., 2015) and forum comments from Hartmann et al. (2019), splitting it by sentences and deriving sentence-level annotations from the original span-level annotations. To do so, we first map the original frame labels to our 9 classes; then we identify spans with matching labels from at least two annotators with a minimum three-word overlap. Following previous methods (Naderi and Hirst, 2017; Hartmann et al., 2019; Park et al., 2022), agreed spans were mapped to the sentence-level: sentences fully covered by agreed spans inherited corresponding frame labels. This results in approximately 63,000 labeled sentences across both genres, but predominantly from the MFC (only 625 sentences from the much smaller Hartmann et al. (2019) dataset). Further details on frame and topic distributions of the training data are provided in Appendix D, Table 7.
Frame Classification Models
To automatically classify frames in our news articles and comments, we fine-tuned three models — RoBERTa-Large (Liu et al., 2019b) (full fine-tuning), LLaMA3.1-8B-Instruct (Grattafiori et al., 2024), and Qwen2.5-7B-Instruct (Qwen et al., 2025) (both parameter-efficient fine-tuning) — using the sentence-level annotated data described above on a random 80/10/10 (training/dev/test) split, stratified by both frame and topic. Fine-tuning parameters are reported in Appendix A, and prompting details in Appendix B.
RoBERTa Llama Qwen
In-Domain 0.66 0.65 0.64
Table 2: Macro-F1 scores of three fine-tuned LMs on MFC + Hartmann data.
X-Domain FrAC (A) FrAC (C)
RoBERTa 0.59 0.77 0.83
Table 3: Macro-F1 scores of Lama cross-domain (X-Domain) and on FrAC articles (A) and comments (C).
Evaluation
Topic SOCC NYT
Articles Comments Avg. C/A Articles Comments Avg. C/A
Gun control 20 855 42 112 28,137 251
Russia-Ukraine 77 2,171 28 310 100,355 323
Trump & Elections 180 7,780 43 247 61,550 249
Healthcare 53 1,636 30 326 67,124 205
Immigration 115 4,394 38 172 33,060 192
LGBT+ Rights 23 803 34 22 3,279 149
Education 90 3,907 43 78 11,614 149
Abortion 17 508 29 38 6,193 163
Israel-Palestine 53 1,794 34 63 7,417 118
Climate Change 311 11,271 36 231 23,990 104
Syria & IS 138 4,195 30 72 3,757 52
Total 1,077 39,614 36.8 1,699 346,476 204
Table 4: SOCC and NYT data statistics by topic and source, including the average number of comments per article, after the pre-processing explained in Section 6.
Table 2 shows little difference in the models’ performance on the 10% MFC/Hartmann test split described above, with RoBERTa-Large slightly outperforming the larger Qwen2.5-Instruct and LLaMA3.1-8B-Instruct. This result aligns with recent literature on frame classification, where BERT-based pre-trained models have consistently demonstrated good performance (Sajwani et al., 2024; Daffara et al., 2025). Since RoBERTa also requires substantially less fine-tuning time (approximately 20 minutes vs. 13 hours for PEFT-tuning of LlaMA and Qwen), all remaining experiments in this paper are based on the fine-tuned RoBERTa model.
To evaluate model generalizability to unseen topics, we train RoBERTa on all but one MFC topics, evaluating on the held-out topic, averaging over all five combinations. Table 3 (X-Domain) revealed an expected, albeit relatively small drop in RoBERTa performance (from 0.66 In-Domain in Table 2 to 0.59), suggesting a reasonable ability to generalize to unseen topics, which is important because the model will be applied to topics not present in the MFC training set.
Finally, we evaluated RoBERTa on the annotated FrAC corpus (Table 3, FrAC), i.e., a data set from different outlets, time points and distinct yet overlapping set of topics (Section 5 and Table 4). The model achieved strong performance on FrAC with an F1 score of 0.77 on articles and 0.83 on comments, demonstrating effective generalization to reader comments despite being trained primarily on formal news articles. This result validates our central methodological contribution: that sentence-level frame classification can successfully bridge the gap between formal news articles and informal reader comments.
Having validated RoBERTa’s performance on our gold-standard FrAC corpus, we then applied it to the full sentence-tokenized SOCC and NYT data sets (Section 4.1) for our large-scale framing effects analysis. [^9] We used the NLTK library https://www.nltk.org/ for sentence tokenization of article and comment texts.
6.2 Dominant Frame Detection
We analyze framing effects between articles and comments in terms of their respective dominant frames. We define the dominant frame as the most frequent sentence-level frame label in a given document. To reduce noise from near ties, we only assign a dominant frame if the most frequent frame meets both of the following criteria: it appears in at least 3 sentences and it represents at least 40% of all sentences in the document. We make an exception for single-sentence texts (comments), whose frame is automatically considered to be dominant. Articles and comments whose most frequent frame fails to meet these combined thresholds were excluded from further analysis.
Evaluation
To assess the reliability of our dominant frame detection heuristics, two authors of the paper examined a sample of 110 articles and comments (55 for each) and assessed if the dominant frame derived by applying the heuristics explained above is correct. The agreement rate between the annotators was 83%. Furthermore, human annotations agreed with the heuristic prediction 88% of the time, showing that our approach effectively captures the dominant frame for both articles (95%) and comments (82%).
6.3 Topic Extraction
We applied BERTopic (Grootendorst, 2022) separately to SOCC and NYT to extract topics from news articles, assigning each article a probability distribution across all identified topics. [^10] We report the full list of parameters used in Appendix C. We analyzed the initial hierarchical topic structures produced by BERTopic and manually merged closely positioned topics with small inter-topic distances and high thematic similarity. The refined topic model was then applied to reader comments. We ultimately assigned each article and each comment the single most likely topic predicted by BERTopic.
This process was conducted independently for both corpora, to avoid any bias in subsequent analyses. The extracted topics showed substantial thematic overlap across data sets, enabling cross-dataset comparison. We manually identified eleven overarching, semantically cohesive topics which we use in our downstream analysis (Table 4).
Evaluation
To assess the quality of topic assignments, we randomly sampled 5 articles and 5 comments from each of the 11 topic groups, yielding a total of 110 texts. Two authors of the paper independently reviewed each assigned topic and judged whether it was acceptable or not. We observed a high degree of inter-annotator agreement (96%). The overall accuracy of the model’s predicted topics, calculated against the labels assigned by the annotators was found to be 93% (on both comments and articles).
6.4 Resulting Data Set
With our framework and validated implementation at hand, we next apply it to analyze framing effects across the SOCC and NYT data sets (Section 4.1) and eleven selected topics (Section 6.3). We only include article-comment pairs where both share the same predicted topic (excluding off-topic comments to study framing effects when discussions remain on-topic) and where dominant frames could be extracted for both articles and comments. We also filter out comments shorter than 5 words, and repeated texts. This final data set consists of nearly 3,000 articles and over 380,000 associated comments, as detailed in Table 4.
7 Framing Effects across Topics and Outlets
Using our corpus of articles and comment frames from two outlets, annotated with both dominant frames and labels, we next address our three research questions (Section 3) in turn.
Figure 2: News article frames are retained in reader comments to varying degrees (X- and Y-axes). The degree of retention correlates strongly across two media outlets: New York Times (NYT) and The Globe and Mail (SOCC).
7.1 To What Extent are Frames Retained vs Changed?
We quantitatively investigated the dominant frames employed in both the news articles and corresponding reader comments across the New York Times (NYT; USA) and The Globe and Mail (SOCC; Canada), analyzing the extent to which the dominant frame of the article was retained in the associated reader comments.
Overall, we find that NYT readers demonstrate stronger frame retention (50.7%) compared to SOCC readers (37.3%). Conversely, at least half of the comments (almost two thirds in SOCC) change away from the dominant article frame indicating that readers often do not retain the original article framing, but change the message. In section 7.3, we analyze these reframing patterns in depth.
Figure 2 displays the frame retention patterns separately for each dominant article frame. Looking at the spread of frames along the x- and y-axis independently, we can see that the extent of retention varies by frame. Specifically, frames such as "Political and Policies", "Economic", and "Health and Safety" show high retention rates. In contrast, "Morality", "Fairness and Equality", and "Cultural Identity" exhibit substantially lower retention rates, suggesting that readers are more likely to reframe content that is presented through frames that require alignment with personal beliefs or values to be accepted (Chong and Druckman, 2007).
[…]
Figure 2 additionally reveals that the reframing patterns are remarkably consistent across NYT and SOCC, i.e., all points cluster close to the diagonal line of perfect correlation. This suggests that our uncovered patterns might reveal a general trend in news reader behavior, rather than being idiosyncratic to a particular outlet or sample, although future analyses across more diverse outlets and countries will need to verify this trend.
Figure 3: Frame retention rates by topic across news outlets. Points above the diagonal line indicate higher retention in NYT, while points below indicate higher retention in SOCC. Topic names are shortened for enhanced visibility.
7.2 Do Framing Effects Vary Across Topics?
Here, we explore the effect of topics on framing effects. We cannot estimate topic effects from the observed data directly, because the effect of topics is confounded by the topic-specific frame distributions. To address this, we estimate the effect of article topic on frame retention using logistic regression mixed effects models. Given a document-response pair $i=(d,r)$:
$y_{i}\sim\text{topic}_{d}+(1|\text{frame}_{d})+(1|\text{id}_{d}),$ (1)
where $y_{i}$ is a binary indicator of frame retention ($\text{frame}_{d}{==}\text{frame}_{r}$). Document (article) topic ($\text{topic}_{d}$) is a fixed effect of topic on frame retention. And $(1|\text{frame}_{d})$ and $(1|\text{id}_{d})$ are fixed effects, controlling for the impact of the article frame ($\text{frame}_{d}$) and article id $(\text{id}_{d})$, respectively. [^11] We control for article ID due to repeated measurements, given that we have multiple comments per article. We then obtain the marginal effects of $\text{topic}_{d}$ on $y$ from the fitted model, removing any influence of the random effects. The resulting probabilities of frame retention per topic are shown in Figure 3 for NYT (y-axis) and SOCC (x-axis).
We find that framing effects vary by topic, but the degree of variance is much higher for NYT compared to SOCC suggesting that these patterns are more outlet specific than the framing-based variation discussed above and shown in Figure 2. This offers an explanation for the inconsistent findings in previous single-topic studies. Some topics exhibit consistent retention patterns across outlets: Coverage of Trump & Elections demonstrates relatively high retention in both data sets (NYT: 31.3%, SOCC: 42.7%), while Healthcare and Education also show moderate to strong retention. International conflict topics are consistently more often reframed by the readers: Israel-Palestine conflict and Syria & IS show the lowest retention rates across both sources, ranging between 18-24%. Other topics reveal outlet-specific differences: Immigration shows substantially higher frame retention in NYT than SOCC, as well as the coverage of the Russia-Ukraine conflict.
7.3 How Do Readers Reframe?
Comments change away from the dominant article frame 50% (NYT) or 63% (SOCC) of the time, so a natural question is how readers reframe articles. Figure 4 summarizes the flow of reframing in the NYT across all topics. Readers predominantly reframe issues into political terms, regardless of the original article framing, but a particularly large proportion of value-laden article frames (“Fairness and Equality”, “Public Opinion”) are recast as “Political and Policies”. [^12] This does not necessarily mean, though, that readers “politicize” all articles – “Political and Policies” frame is also the most reframed one, and its relative proportion in comments is actually lower than in the articles. This pattern, albeit to a lesser degree, also emerges in SOCC (Figure 6 in Appendix), where readers also often recast different frames into “Economy”.
We also investigated topic-specific patterns of reframing for topics that included at least 100 articles. For instance, in the case of Immigration (Table 5), SOCC readers both reframe “Cultural Identity” into “Political and Policies” terms, and vice versa, suggesting a tight interconnection of these two perspectives. NYT readers, on the other hand, primarily reframe into ‘harder’ perspectives like “Economic” and “Legality and Crime” and interestingly, among these top five reframings (as shown in the table), the “Economic” framing occurs only on the reader side. Table 9 in the Appendix shows the analysis for three additional topics, revealing varying trends across topics and outlets.
Figure 4: Reframing patterns in the New York Times showing the flow from primary article frames (left) to primary comment frames (right) with proportion of frames in articles and comments, respectively. We plot standardized frame proportions in articles, to reduce visual clutter.
Reframing Pattern N
SOCC Cultural Identity → Political & Policies 134
Political & Policies → Cultural Identity 131
Health & Safety → Cultural Identity 120
Economic → Political & Policies 108
Health & Safety → Political & Policies 99
NYT Legality & Crime → Political & Policies 1388
Political & Policies → Economic 1311
Political & Policies → Legality & Crime 1139
Legality & Crime → Economic 1024
Political & Policies → Cultural Identity 950
Table 5: Five most common reframings of the topic ‘Immigration’ in SOCC (top) and NYT (bottom) as ‘dominant article frame $\rightarrow$ dominant comment frame’.
8 Discussion
Unlike previous computational approaches to framing detection and analysis, our study is centered on a constructionist understanding of framing as an interactive process where readers selectively engage, modify, and reconstruct media frames based on their own interpretative resources  (Neuman et al., 1992; Scheufele, 1999; Gorp, 2007). While existing computational methods treat framing as isolated classification tasks focused solely on source content, we examined the dynamic interaction between media framing and audience reception. In uncovering systematicities (RQ1; Figure 4) and topic-specific differences (RQs 2-3) in a large-scale study, we bridged disconnect between large-scale yet isolated NLP approaches and small-scale communication studies that lack generalizability across topics and contexts. In applying one framing conceptualization across eleven topics, our findings suggest that inconsistent results in prior work stem from topic- and outlet-specific variation in framing effects. We also found consistent trends, such as value-laden frames (“Fairness and Equality”, “Morality”) being more likely to be reframed. This is in line with theoretical formulations of framing, according to which individuals more readily accept frames that match their pre-existing values (Chong and Druckman, 2007).
Our analysis reveals outlet-specific trends — e.g., NYT readers’ political reframing and SOCC readers’ economic focus — likely reflect interactions between editorial stance, audience demographics, and national context. While frame-based patterns are relatively stable across outlets, topic effects vary more substantially. Since our comparison conflates country (US vs. Canada), editorial leaning (center-left vs. center-right), and time period (2012–2016 vs. 2017–2018), we refrain from strong interpretation. Future work should disentangle these factors to clarify framing dynamics and inform broader interdisciplinary research.
A key methodological contribution of our framework lies in demonstrating reliable frame classification across fundamentally different text genres – from formal news articles to informal reader comments. This cross-genre generalization addresses a critical gap in computational framing analysis, where models typically operate within homogeneous text types. Our validation shows that sentence-level frame classification can successfully detect frames in structurally and stylistically different content, with strong performance on our human-annotated data set FrAC.
Our study builds on the MFC, a widely-used framework in NLP. Yet, it is by no means the only or the optimal framework, and it has been criticized for the broad nature and lack of theoretical grounding of its frames (Ali and Hassan, 2022). We note that our method is framework-agnostic: alternative framing theories could be readily integrated, whether more generic frameworks (e.g., Semetko and Valkenburg (2000)), or issue-specific taxonomies tailored to particular domains, a rich direction for future work. Similarly, our methods for dominant frame detection and sentence-level annotation, while supporting discovery of major trends, limit finer-grained framing analyses, which could be explored in future work.
Finally, our framework holds potential relevance for other NLP tasks including argumentation interpretation and generation, and even disinformation detection. The systematic patterns we identify in how audiences reconstruct frames could inform models of persuasive text generation, helping systems understand how different framings might be received and transformed by target audiences.
9 Conclusion
We introduced a scalable computational framework for analyzing media framing effects as an interactive process between source content and audience responses, grounded in a constructionist view on framing. Through analysis of over 3,000 news articles and 380,000 comments across two outlets and eleven topics, we demonstrate that readers reframe content at least as frequently as they retain original article frames. Frame retention varies systematically by frame type, with value-laden frames being more likely to be reframed, and when reframing occurs, readers consistently shift toward political interpretation regardless of the original framing. To support our analysis, we contributed a robust framing scheme from an established taxonomy, together with robust classifiers that work across different document types, providing a foundation for future research on media framing effects.
Acknowledgment
This paper was written with the support from the Melbourne Research Scholarship by the University of Melbourne provided to MG. This work was supported by the Australian Research Council Discovery Early Career Research Award (Grant No. DE230100761). We thank Damian Curran, Thomas Hollow, Jemima Kang, Tom Monaghan, Gordon William Smith, and Katie Warburton for their annotations.
Appendix A Frame Classifiers
We report the parameters used for fine-tuning the two LLMs — Llama and Qwen — and RoBERTa for the frame classification task. The LLMs were fine-tuned using Unsloth [^13] https://unsloth.ai/, applying parameter-efficient fine-tuning via Low-Rank Adaptation (LoRA) (Hu et al., 2022), while RoBERTa was fully fine-tuned. All training was conducted on high-performance computing infrastructure with NVIDIA A100 GPUs. We also include the final prompt used for the task in Table 6.
A.1 LLaMA3-8B-Instruct and Qwen2.5-Instruct
We used the same set of parameters on Unsloth for both models.
  • Max Sequence Length: 3000 tokens
  • Batch Size: 8 (gradient accumulation = 8, effective batch = 64)
  • Number of Epochs: 6
  • Learning Rate: 2e-4
  • LoRA Configuration: r = 16, $\alpha$ = 16, dropout = 0, target modules = [q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, down_proj]
  • Precision: bf16 (if supported) or fp16
  • Optimizer: AdamW (8-bit)
  • Scheduler: Linear with 10% warm-up ratio
  • Regularization: Weight decay = 0.01
  • Early Stopping: Patience = 3 epochs
  • Checkpointing: Save best model by validation loss at end of each epoch
  • Packing: Enabled
A.2 RoBERTa
  • Maximum sequence length: 70 tokens
  • Data splitting:
    • Training set: 80%
    • Validation set: 10%
    • Test set: 10%
  • Batch size: 64
  • Learning rate: 5e-6
  • Number of epochs: 10
  • Weight decay: 0.01
  • Optimizer scheduler: Linear learning rate scheduler with 5% warmup
  • Early stopping: Patience of 2 epochs based on validation F1 score
  • Evaluation and model saving: After every epoch, best model saved based on F1 score
  • Loss function: Cross-entropy
Appendix B LLMs Prompt
We experimented with various prompting strategies to balance performance and prompt length. We found that including an elaborated description of the category definitions from Card et al. (2015), along with clear task instructions, consistently yielded the best results. The final prompt is reported in Table 6.
Prompt Used for Fine-Tuning
You are an expert in discourse and framing analysis. You are analyzing how texts are framed in their coverage of topics. Your task is to identify the frames used in the text into one of nine frame categories based on how the issue is presented and which aspects are emphasized.
Instructions: 1. Read the text carefully. 2. Identify the frame that shapes the article’s narrative. Only ONE frame applies. 3. Output ONLY the NUMERICAL code (e.g. "2") for those frames first, followed by a new line. Do not include any other text or explanation. 1 - Economic: The costs, benefits, or financial implications of the issue (individual, community, or economy)
2 - Morality: Perspectives compelled by religion, ethics, or sense of social responsibility
3 - Fairness and Equality: Equality or inequality in how laws and resources are applied or distributed
4 - Legality and Crime: Laws, judicial interpretations, crime rates, legal consequences
5 - Health and Safety: Healthcare, disease, sanitation, mental health, infrastructure safety
6 - Cultural Identity: Social norms, values, customs, and cultural preservation
7 - Public Opinion: Public attitudes, polling, consequences of diverging from popular opinion
8 - Security and Defense: Protection, national security, border security, social stability
9 - Political and Policies: Political discourse, partisanship, lobbying, legislative processes
10 - Other: If the sentence is not framed in any of the proposed categories, select this label
Table 6: Prompt template format used for fine-tuning the models.
Appendix C BERTopic Configuration
We trained a BERTopic model (Grootendorst, 2022) for clustering textual topics with the following configuration:
  • Embedding Model: all-mpnet-base-v2 from the SentenceTransformers library.
  • Vectorizer: CountVectorizer with the following parameters:
    • stop_words: English stopwords from NLTK
    • min_df: 5
    • ngram_range: (1, 2)
  • UMAP Dimensionality Reduction:
    • n_neighbors: 15
    • n_components: 5
    • min_dist: 0.0
    • metric: cosine
    • random_state: 42
  • Clustering Model: HDBSCAN with:
    • min_cluster_size: 15
    • min_samples: 7
    • metric: euclidean
    • cluster_selection_method: eom
    • prediction_data: True
  • BERTopic Parameters:
    • calculate_probabilities: True
    • verbose: True
    • n_gram_range: (1, 2)
    • min_topic_size: 5
Appendix D Processed Training Data
We report the statistics of the final processed data set, as explained in Section 6.1. Table 7 reports the number of final samples per label across topics, from both the MFC and Hartmann et al. (2019).
Frame Guns Death Penalty Immig. Same-Sex Tobacco Hartmann Total
Legality and Crime 4,942 9,215 4,111 3,306 1,376 344 23,294
Political and Policies 7,405 1,006 3,436 3,357 1,853 211 17,268
Security and Defense 947 163 770 11 - - 1,891
Health and Safety 938 740 534 158 1,409 - 3,779
Economic 810 156 1,294 441 2,303 70 5,074
Morality 183 573 147 999 64 - 1,966
Cultural Identity 636 432 722 552 459 - 2,801
Public Opinion 1,089 161 470 967 130 - 2,817
Fairness and Equality 60 677 284 681 34 - 1,736
Other 275 451 253 1,716 305 - 3,000
Total 17,285 13,574 12,021 12,188 7,933 625 63,626
Table 7: Distribution of frame categories across issue topics in the training data used for fine-tuning the frame classification models. The data combines the Media Frames Corpus with additional annotations from Hartmann et al. (2019).
Appendix E Annotation Instructions and Interface
This section includes the instructions which were presented to each annotator on the web interface developed for the task (a screenshot of the interface is reported in Figure 5).
Figure 5: An example from the annotation task on the deployed interface.
Understanding the Influence of News Framing on Readers
Hi there! Thanks for agreeing to participate in this annotation task.
If you’re viewing this on a phone or tablet, we recommend switching to a laptop for better comfort and readability.
In this project, we’re investigating the different formulations of news in terms of frames.
What Are Frames?
Frames are the perspectives or angles through which information is presented in media. They highlight certain aspects of an issue while downplaying others, influencing how audiences understand and respond to news.
Think of frames as different “lenses” that focus attention on specific elements of a story while leaving others in the background.
Frames answer the question: “How is the issue at hand being presented?”
Your Task
Your task is to identify which frames are used in short snippets of text from various sources. This helps us understand how different issues are presented and framed in different contexts.
How the Annotation Works
You will see one text at a time. Please read each text carefully and choose the frame that best represents how the information is being conveyed by clicking on the corresponding label (colourful boxes).
If you think more than one frame applies, please select all that you believe are relevant.
If none of the provided categories apply, please select the label “Other.”
Please note: the examples all focus on same-sex marriage for clarity. The real annotation task covers a variety of different issues.
Once you start annotating, you can click on the “Instructions” button at any time to retrieve this information.
Framing Categories
<Displays list of framing categories as in Table 8>
Frame Description Example
Economic The costs, benefits, or financial implications of the issue (individual, community, or economy). Legalizing same-sex marriage would boost state economies through increased spending on weddings and tourism.
Morality Perspectives compelled by religion, ethics, or sense of social responsibility. Same-sex marriage is morally wrong and goes against the teachings of our faith.
Fairness and Equality Equality or inequality in how laws and resources are applied or distributed. Denying same-sex couples the right to marry is a clear violation of equal protection under the law.
Legality and Crime Laws, judicial interpretations, crime rates, legal consequences. The Supreme Court ruled that bans on same-sex marriage are unconstitutional and violate civil liberties.
Health and Safety Healthcare, disease, sanitation, mental health, infrastructure safety. Allowing same-sex marriage is an endorsement of unsafe sexual practices and will contribute to the spread of disease.
Cultural Identity Social norms, values, customs, and cultural preservation. Same-sex marriage threatens the traditional concept of family that our culture has upheld for generations.
Public Opinion Public attitudes, polling, consequences of diverging from popular opinion. A recent poll shows that 65% of Americans now support same-sex marriage, up from 30% a decade ago.
Security and Defense Protection, national security, border security, social stability. Same-sex marriage could undermine traditional social structures and thus contribute to community stability and security.
Political and Policies Political discourse, partisanship, lobbying, legislative processes. The governor vetoed a bill legalizing same-sex marriage, citing pressure from conservative lawmakers.
Other If the sentence is not framed in any of the proposed categories, select this label.
Table 8: Framing Categories used in this work.
Practice Phase
Before starting the actual annotation task, we’ll practice together with three examples.
During the practice phase:
  • You will be shown one sentence at a time.
  • You should select the appropriate frame label(s) by clicking on the colored boxes.
  • After you make your selection, the correct answer will appear underneath with an explanation.
Take your time to read each sentence carefully and consider which perspective or angle is being emphasized.
You can consult the examples you will see during the practice at any time during the annotation by clicking on the “Tutorial Example” button on the top-right.
Final Notes
If you have any questions, please contact the organiser.
For more information regarding the framing labels and theory, please refer to the Policy Frames Codebook and this paper.
Happy annotating!
Appendix F Reframing Patterns in SOCC
Figure 6: Reframing patterns in the SFU Opinion Corpus (SOCC) showing the flow from primary article frames (left) to primary comment frames (right). We print the original % of frames in articles and comments, respectively. This is reflected on the bar height on the right, but not on the left to reduce visual clutter.
Appendix G Reframing Patterns by Topic
Topic SOCC NYT
Reframing Pattern Count Reframing Pattern Count
Climate Change Political and Policies → Economic 1126 Political and Policies → Economic 1590
Economic → Political and Policies 907 Economic → Political and Policies 857
Political and Policies → Cultural Identity 185 Political and Policies → Morality 690
Political and Policies → Public Opinion 181 Political and Policies → Health and Safety 686
Economic → Cultural Identity 156 Political and Policies → Cultural Identity 465
Immigration Cultural Identity → Political and Policies 134 Legality and Crime → Political and Policies 1388
Political and Policies → Cultural Identity 131 Political and Policies → Economic 1311
Health and Safety → Cultural Identity 120 Political and Policies → Legality and Crime 1139
Economic → Political and Policies 108 Legality and Crime → Economic 1024
Health and Safety → Political and Policies 99 Political and Policies → Cultural Identity 950
Syria & IS Security and Defense → Political and Policies 405 Security and Defense → Political and Policies 321
Political and Policies → Security and Defense 360 Security and Defense → Morality 165
Security and Defense → Morality 127 Health and Safety → Political and Policies 149
Security and Defense → Cultural Identity 120 Health and Safety → Morality 123
Political and Policies → Morality 96 Security and Defense → Cultural Identity 106
Trump & Elections Political and Policies → Cultural Identity 407 Political and Policies → Morality 2516
Political and Policies → Public Opinion 293 Political and Policies → Cultural Identity 2373
Political and Policies → Economic 206 Political and Policies → Health and Safety 1362
Political and Policies → Fairness and Equality 178 Cultural Identity → Political and Policies 1145
Cultural Identity → Political and Policies 145 Political and Policies → Economic 1039
Table 9: Top 5 reframing patterns by topic for SOCC and NYT outlets, showing the transition from one frame to another and the frequency count.

### v2

@@
 When a news article describes immigration as an "economic burden" or a "humanitarian crisis," it selectively emphasizes certain aspects of the issue.
-Although framing shapes how the public interprets such issues, audiences do not absorb frames passively but actively reorganize the presented information.
+Although this framing shapes how the public interprets such issues, audiences do not absorb frames passively but actively reorganize the presented information.
 While this relationship between source content and audience response is well-documented in the social sciences, NLP approaches often ignore it, detecting frames in articles and responses in isolation.
@@
 Methodologically, we refine frame labels and develop a framework that reconstructs dominant frames in articles and comments from sentence-level predictions, and aligns articles with topically relevant comments.
-Applying our framework across eleven topics and two news outlets, we find that frame reuse in comments correlates highly across outlets, while topic-specific patterns vary.
+Applying our framework across eleven topics and two news outlets, we find that frame reuse in comments correlates highly across outlets, and that readers often selectively engage with frames of the articles.
 We release a frame classifier that performs well on both articles and comments, a dataset of article and comment sentences manually labeled for frames, and a large-scale dataset of articles and comments with predicted frame labels.
-[^1] We released our at: https://anonymous.4open.science/r/FrAC-A8E3/; and fine-tuned frame classifier: https://huggingface.co/mattdr/sentence-frame-classifier
+[^1] We released our data at: https://github.com/mattguida/FrAC; and fine-tuned frame classifier: https://huggingface.co/mattdr/sentence-frame-classifier
 1 Introduction
@@
 However, audiences do not passively absorb frames but rather "actively filter, sort and reorganize information in personally meaningful ways in the process of constructing an understanding of public issues" (Neuman et al., 1992, pg.77).
-Such a constructionist understanding emphasizes that frames cannot be fully understood without considering the interaction between source and receiver (Scheufele, 1999; Gorp, 2007; Hubner and Dixon, 2023).
-This is particularly evident on news websites with comment threads that create immediate opportunities for readers to publicly respond to framed content — a paradigm shift from one-way communication to participatory discourse where audiences actively engage with and reconstruct news content (Risch and Krestel, 2020).
+The constructionist understanding of framing emphasizes that frames cannot be fully understood without considering the interaction between source and receiver (Scheufele, 1999; Gorp, 2007; Hubner and Dixon, 2023), as part of an interactive and dynamic process in which frames activate and reasonate (or fail to) with personal schemas through active meaning construction and intepretation.
+This is particularly evident on news websites with comment threads that create immediate opportunities for readers to publicly respond to framed content – a paradigm shift from one-way communication to participatory discourse where audiences actively engage with and reconstruct news content (Risch and Krestel, 2020).
 However, current research has yet to fully explore the relationship between media framing and audience reception.
@@
 Figure 1: Overview of our framework. Bottom: Article with two comments illustrating reframing (Comment 1; changes the article frame) and frame retention (Comment 2; keeps the article frame).
+If framing is truly a process in which frames are actively constructed and interpreted by an audience (Scheufele, 1999; Gorp, 2007; Hubner and Dixon, 2023), then the audience’s response is not merely a posited outcome, but a fundamental part of the analysis that has been largely overlooked or observed at a limited scale.
 To address this gap, we propose a computational framework that investigates at scale the relationship between how information is framed in news articles and how readers respond to such framed content in associated comments.
-Drawing from the constructionist understanding of framing (Gorp, 2007) and using an established framing taxonomy (Boydstun et al., 2014; Card et al., 2015), we examine both source content and audience responses as interconnected elements of a broader communicative process (Hubner and Dixon, 2023).
-Throughout this paper, we use media framing effects to describe the relation between news article framing and that of reader comments, as an umbrella term that covers two phenomena: frame retention occurs when readers adopt the dominant frame used in the source article in their comments, while reframing occurs when readers shift to a different dominant frame in their comments.
-Figure 1 illustrates both, with an example.
-Our large-scale analysis across eleven diverse topics and two news outlets [^2] The New York Times (USA) and The Globe and Mail (Canada) reveals systematic patterns of media framing effects.
-We find that (a) frame retention rates vary by source document frame, and this variation correlates highly across the two outlets, with value-laden frames being more often reframed; (b) retention patterns vary systematically by topic, and (c) when reframing, readers predominantly shift discussions toward political interpretation regardless of the original article frame.
+Using an established framing taxonomy (Boydstun et al., 2014; Card et al., 2015), we examine both source content and audience responses as interconnected elements of a broader communicative process (Hubner and Dixon, 2023).
+Throughout this paper, we use the phrase media framing effects to describe the relation between news article framing and that of reader comments, as an umbrella term that covers two phenomena.
+Frame retention occurs when readers adopt the dominant frame [^2] We define the dominant frame as the one that covers the most sentences of an article, but at least 40% (see Section 6.2).
+used in the source article in their comments.
+Reframing occurs when readers shift to a different frame in their comments: either a minor frame in the article (selective reframing), or a completely new, unmentioned frame (complete reframing).
+Figure 1 illustrates this with an example.
+Our large-scale analysis across eleven diverse topics and two news outlets (The New York Times and The Globe and Mail) reveals systematic patterns of media framing effects.
+We find that (a) frame retention rates vary by source document frame, and this variation correlates highly across the two outlets, with value-laden frames being more often reframed; (b) retention patterns vary systematically by topic, and (c) audiences engage in selective reframing, adopting secondary frames rather than completely accepting or rejecting primary frames.
 We make three main contributions: (1) a computational framework for investigating the relationship between source frames (articles) and frames in associated responses (comments), (2) a frame classification model that effectively generalizes across different text types (news articles and reader comments) and topics, and (3) two data sets: a manually labeled evaluation corpus (FrAC) and a large-scale data set of articles and comments aligned by topic group and labeled with dominant frames, for reproducibility and supporting future research.
@@
 Cognitive Studies on Framing Effects
-Cognitive research established that news frames significantly influence audience interpretation using controlled experimental designs, where participants were assigned to different framing conditions of the same issue and asked to report their thoughts.
+Cognitive research has established that news frames significantly influence audience interpretation using controlled experimental designs, where participants were assigned to different framing conditions of the same issue and asked to report their thoughts.
 Valkenburg et al.
@@
 (1997) showed that frames not only guide interpretation but also stimulate new thoughts and feelings not explicitly present in the original content.
-Nelson and Oxley (1999) and Shen (2004) showed that framing primarily influences the importance of existing beliefs rather than changing the beliefs themselves, with frames most effective when resonating with pre-existing schemas.
-While producing valuable insights, these studies were restricted to small-scale experimental settings and narrow demographic representation.
-This limited their generalizability to framing effects in natural media environments with direct audience responses.
+Nelson and Oxley (1999) and Shen (2004) showed that framing primarily influences the importance of existing beliefs rather than changing them, with frames being the most effective when they resonate with pre-existing schemas.
+These studies were restricted to small-scale experimental settings and narrow demographic representation, and thus could not generalize to natural media environments with direct audience responses.
 Media Framing Effects in Communication Studies
@@
 Holton et al.
-(2014) analyzed online health coverage across major outlets and found no evidence of frame retention in comment threads, noting a disconnect between thematic news frames and predominantly episodic comment frames.
+(2014) analyzed online health coverage across major outlets and found no evidence of frame retention in comment threads.
 Muñiz et al.
@@
 Computational Approaches
-NLP approaches to framing have evolved from exploratory topic-modeling (DiMaggio et al., 2013) to structured classification tasks (for a review, see Ali and Hassan (2022)).
+NLP approaches to framing have evolved from exploratory topic-modeling (DiMaggio et al., 2013) to classification tasks (for a review, see Ali and Hassan (2022)).
 The Media Frame Corpus (MFC, Card et al.
 (2015)), based on Boydstun et al.
-(2014), has been particularly influential, spurring various frame classification methods including probabilistic approaches (e.g.
-Johnson and Goldwasser (2018)), neural networks (e.g.
-Naderi and Hirst (2017)), and pre-trained language models (e.g.
-Khanehzar et al.
+(2014), has been particularly influential, spurring various frame classification methods including probabilistic approaches (e.g.,Johnson and Goldwasser (2018)), neural networks (e.g., Naderi and Hirst (2017)), and pre-trained language models (e.g., Khanehzar et al.
 (2021)).
@@
 (2025) tested cross-linguistic transfer using both fine-tuned pre-trained language models and zero-shot LLM classification on Portuguese news articles.
-However, these studies either focused on cross-linguistic transfer, or on the diagnosis of the problem of cross-topic and genre generalization but fell short of offering a solution to the problem, which we do here.
+However, these studies either focused on cross-linguistic transfer or on the diagnosis of the problem of cross-topic and genre generalization.
+They fell short of offering a solution to the problem, which we do here.
 3 Research Questions
-The gaps in existing research motivate three fundamental questions about media framing effects in readers comments:
-  1. 1.
-    To what extent are frames retained vs changed?
-Previous studies have shown mixed results.
-We investigate at scale whether article frames consistently influence comment frames, and whether this influence varies systematically across different types of frames.
-  2. 2.
-    How do framing effects vary across topics?
-Topic-specific studies cannot reveal whether framing effects represent universal phenomena or topic-specific patterns.
-We examine whether framing effects are consistent or vary based on the topic being discussed.
-  3. 3.
-    When readers do not retain the dominant frame of the article, how do they reframe it?
-Understanding reframing patterns is crucial for comprehending how audiences actively reconstruct media messages.
-We examine systematic tendencies in how readers shift framing when they reject source frames.
+Our goal is to characterize and quantify the constructivist view on framing at scale.
+We formulate research questions that have been raised in the social sciences, albeit on a small scale often regarding a single topic and a specific framing inventory which makes it difficult to generalize results across scenarios.
+With orders of magnitude more data that covers a range of topics and is analyzed with a unified framing inventory, we are well positioned to answer the following research questions.
+RQ1: To what extent are dominant article frames retained in reader comments?
+We first focus on the extent to which the most prominent frame in the article is reflected in the reader comments.
+From a constructivist point of view, this assesses the extent to which a chosen article frame sticks with the readership.
+Next, we address the fact that prior work was often topic-specific and led to inconclusive results about the extent of frame retention.
+Here, we can quantify the effect of topics on frame retention within a unified framework.
+We ask
+RQ2: How does frame retention vary across topics?
+Finally, we examine how readers how audiences reconstruct media messages when they do reframe (i.e., they choose not to retain the dominant article frame).
+Our framework allows us to paint a more nuanced picture on reframing than previous work.
+First, in line with prior studies we quantify how often readers adopt a frame that was never used in the article (complete reframing).
+In addition, we also explore how often readers shift from the dominant article frame to a secondary article frame (selective reframing).
+This leads to our final research question:
+RQ3: If readers reject the dominant article frame, how do they reframe?
+We next introduce our framework and methodology, before we address our research question in Section 7.
 4 Framework
@@
 4.1 Data
-Our framework requires a data set of source documents $d_{i}$ (e.g.
-news articles, headlines, tweets, etc.) each paired with a set of response documents $R_{d_{i}}=\{r_{1},r_{2},\ldots\}$ (e.g.
-comments, retweets, blog posts, etc.) directly linked to the source ($D=\{(d_{i},R_{d_{i}})\}_{i=1}^{N}$).
-For the purposes of this study, we use two existing data sets of news articles with their associated comments:
+Our framework requires a data set of source documents $d_{i}$ (e.g., news articles, headlines, tweets, etc.) each paired with a set of response documents $R_{d_{i}}=\{r_{1},r_{2},\ldots\}$ (e.g., comments, retweets, blog posts, etc.) directly linked to the source ($D=\{(d_{i},R_{d_{i}})\}_{i=1}^{N}$).
+For the purpose of this study, we use two existing data sets of news articles with their associated comments:
   • SOCC (The SFU Opinion Comments Corpus; Kolhatkar et al.
@@
 with over two million comments associated with approximately 9,000 articles published between January-May 2017 and January-April 2018.
-We note that the framework generalizes to any corpus where source documents are mapped to audience responses.
+This pairing of source documents with response documents directly operationalizes the constructionist understanding of framing.
+In our analysis, news articles represent the ’source framing’ – how journalists and media outlets choose to present issues through particular frames.
+Reader comments represent the ’audience response’ – how readers actively interpret, accept, modify, or reject those frames based on their own cognitive schemas and values.
+We focus exclusively on top-level comments that directly respond to articles, excluding nested replies within comment threads, as we want to capture direct audience responses to the original framed content.
 4.2 Framework Structure
-The framework consists of three core components that work in concert to analyze framing dynamics between articles and comments (Figure 1):
+The framework consists of three core components that work in concert to analyze framing dynamics between articles and comments (Figure˜1):
   1. 1.
@@
   2. 2.
-    Dominant Frame Detection: In line with prior work in cognitive and communication studies (Section 2), we identify and compare a single, dominant frame (defined as the most recurring one; Section 6.2) in source documents and responses.
+    Dominant Frame Detection: In line with prior work in cognitive and communication studies (Section 2), we identify and compare a single, dominant frame (defined as the most recurring one) in source documents and responses.
 We thus require a method to determine the dominant frame in each input document.
+We define the dominant frame as the most frequently occurring sentence-level frame within a document (see Section 6.2 for details).
+In doing so, we can distinguish the retention of the primary article frame from selective reframing into a secondary frame of the original article.
   3. 3.
     Topic Identification: We assign each source document and reader response to a topic, for two reasons.
-First, assigning topics to articles allows us to perform analysis of framing effects over multiple topics, advancing over previous, topic-specific studies (Section 2) and potentially explaining their divergent results.
+First, assigning topics to articles allows us to perform analysis of framing effects over multiple topics, advancing over previous, topic-specific studies (Section˜2) and potentially explaining their divergent results.
 Second, we want to capture framing effects only for responses where commentators stay on topic with the source, rather than digress into a different topic (with potentially a different frame, which can lead to underestimation of retention rates).
-Below, we first introduce a new corpus that we annotated for the frame classification subtask (Section 5).
-Section 6 explains and evaluates our implementations of the three framework components introduced above.
-Finally, we apply the framework to construct a data set (Section 6.4) that we use to analyze media framing effects (Section 7).
+Below, we first introduce a new corpus that we annotated for the frame classification subtask (Section˜5).
+Section˜6 explains and evaluates our implementations of the three framework components introduced above.
+Finally, we apply the framework to construct a data set (Section˜6.4) that we use to analyze media framing effects (Section˜7).
 5 The FrAC Corpus
@@
 5.1 Frame Labels and Granularity
-We devise a robust frame label set, building on the Media Frames Corpus (MFC).
-The MFC is a foundational data set in computational framing analysis based on Boydstun et al.
-(2014) theoretical framework operationalizing framing through 15 topic-agnostic frame categories.
+We devise a robust frame label set, building on the Media Frames Corpus (MFC), a foundational data set in computational framing analysis based on the Boydstun et al.
+(2014)’s theoretical framework operationalizing framing through 15 generic frame categories.
 The original MFC corpus Card et al.
-(2015) consists only of news articles; however, more recently the framework was successfully applied to other types of documents, including online forum comments Hartmann et al.
+(2015) consists only of news articles; however, more recently the framework has been successfully applied to other types of documents, including online forum comments Hartmann et al.
 (2019), suggesting suitability for our task that involves both long, formal documents and short, less formal responses.
-However, it has been noted (e.g., Ali and Hassan (2022); Daffara et al.
+First, to apply the MFC labels across articles and comments, we need to define a granularity that is meaningful for both of them.
+While most NLP studies on the MFC modeled article-level dominant frames, this is not feasible for reader responses which are often brief and composed of single sentences.
+We therefore chose the sentence level – the smallest shared textual unit between the two types of text – as our basic unit of analysis.
+[^5] Section 6.2 explains how we derive a dominant document frame from sentence-level frames.
+Second, we improve the robustness of the MFC label set as it has been found (e.g., Ali and Hassan (2022); Daffara et al.
 (2025)) that some MFC frames are poorly defined and difficult to distinguish, leading to low inter-annotator agreement and model performance.
-First, to apply the MFC labels across our two text types (articles and comments), we need to define a granularity that is meaningful across both of them.
-While most of NLP studies on MFC looked at article-level dominant frames in the MFC, it is not feasible for comments (or other response documents), which are often brief and composed of single sentences.
-We therefore chose to use the sentence level – the smallest shared textual unit between the two types of text.
-[^5] Section 6.2 explains how we derive a dominant document frame from sentence-level frames.
-Second, to improve robustness, we isolated highly confused frame labels in the original MFC data set.
-To do so, we first identified spans-level MFC annotations (of ${\geq}3$ words) to which two or more annotators assigned different frame labels.
+To do so, we isolated highly confused frame labels in the original MFC data set by first identifying span-level MFC annotations (of ${\geq}3$ words) to which two or more annotators assigned different frame labels.
 We examined such inconsistent annotations using confusion matrices – both for all MFC data and separately for each topic included in the MFC corpus.
@@
 In this case, we conclude that the confusion is due to the topic specifics rather than meaning of the label, and keep such semantically distinct labels separate.
-resulting in two merges of two frames (Table 1).
+resulting in two merges of two frames (Table˜1).
 To further improve robustness of the label set, we removed the two least frequent classes (“External regulation and reputation” and “Capacity and resources”) which accounted for less than 0.7% of the data.
 We finally obtained 9 frame categories, plus a generic “Other” frame to account for instances that do not match any of our nine frames.
-Table 1 lists the frames, with full definitions in Appendix Table 8.
-1) Legality and Crime, 2) Political and Policies+, 3) Economic, 4) Health and Safety, 5) Cultural Identity, 6) Public Opinion, 7) Morality, 8) Fairness and Equality, 9) Security and Defense, 10) Other.
-Table 1: The FrAC frame label set.
-Merged MFC labels "Legality, constitutionality and jurisprudence" and "Crime and punishment".
-+ Merged MFC labels "Policy prescription and evaluation" and "Political".
+This category captures both genuinely unframed content and framed content that falls outside our nine established categories.
+Table 1 lists the frames, with the definitions following those of 10, as depicted in the Appendix by Table 10.
+1) Legality and Crime, 2) Political and Policies+,
+3) Economic, 4) Health and Safety, 5) Cultural Identity, 6) Public Opinion, 7) Morality, 8) Fairness and Equality, 9) Security and Defense, 10) Other.
+Table 1: The FrAC frame label set. Label mergers: $\{$"Legality, constitutionality and jurisprudence", "Crime and punishment"}\}; ${=}\{$"Policy prescription and evaluation", "Political"}\}.
 5.2 Human Annotation
-To validate our frame label set (and model predictions) we collect human annotation on a subset of the SOCC corpus (Section 4.1).
+To validate our frame label set (and model predictions) we collected human annotation on a subset of the SOCC corpus (Section˜4.1).
 Annotation Setup
@@
 [^8] This study was approved by Human Ethics Committee (Referance No.
-2025-32561-67793-4), and has been taken out to the according ethical standards.
+2025-32561-67793-4) and has been taken out to the according ethical standards.
 Annotators were compensated with vouchers over 50 USD, well above local minimum wage requirements.
@@
 Inter-Annotator Agreement
-As annotators could assign multiple labels to each sentence, we assessed inter-annotator agreement in two ways.
+Given our multi-label setup, we assessed inter-annotator agreement in two ways.
 Following prior work Arora et al.
 (2025), we use label-wise Krippendorff’s $\alpha$ to assess per-label agreement, and the Jaccard index to measure the averaged overlap in label sets between annotators.
-The averaged Krippendorff’s $\alpha$ across all labels and annotators indicates moderate to high agreement ($\alpha=0.76$), with some variance across frame labels (min for Morality with $\alpha=0.63$, max for Economic with $\alpha=0.89$).
-However even the category with lowest $\alpha$ reaches moderate agreement.
-The Jaccard index of 0.79 confirms this trend, suggesting a substantial overlap across label sets.
+The averaged Krippendorff’s $\alpha$ across all labels and annotators indicates moderate to high agreement ($\alpha=0.76$), with some variance across frame labels (Table 2), however even the frame category with the lowest $\alpha$ reaches moderate agreement.
+We also compute an averaged Jaccard index of 0.79 which confirms this trend, suggesting a substantial overlap across label sets.
 This shows that a more robust set of labels applied at a sentence level allows to achieve substantial agreement despite the purported subjectivity of the task.
+Label $\alpha$ Label $\alpha$
+Economic 0.89 Political / Policy 0.78
+Health / Safety 0.87 Cultural Identity 0.76
+Legality / Crime 0.82 Other 0.66
+Security / Defense 0.79 Fairness / Equality 0.65
+Public Opinion 0.79 Morality 0.63
+Table 2: Krippendorff’s $\alpha$ per label, averaged across annotation sessions and sorted in descending order.
 Gold Standard
@@
 We kept only the cases where both adjudicators agreed with one of the original annotations, discarding 13 sentences.
-This process resulted in a validation data set of 556 sentences, which we use in Section 6.1 to assess model performance.
+This process resulted in a validation data set of 556 sentences, which we use in Section˜6.1 to assess model performance.
 6 Methodology
@@
 We train frame classifiers that predict sentence-level media frames for both comments and articles, following the framing schema defined in Section 5.1.
-To do so, we require suitable training data from both genres, and a robust evaluation setup.
-Frame Classification Data
-We source articles from the Media Frames Corpus (Card et al., 2015) and forum comments from Hartmann et al.
-(2019), splitting it by sentences and deriving sentence-level annotations from the original span-level annotations.
-To do so, we first map the original frame labels to our 9 classes; then we identify spans with matching labels from at least two annotators with a minimum three-word overlap.
-Following previous methods (Naderi and Hirst, 2017; Hartmann et al., 2019; Park et al., 2022), agreed spans were mapped to the sentence-level: sentences fully covered by agreed spans inherited corresponding frame labels.
-This results in approximately 63,000 labeled sentences across both genres, but predominantly from the MFC (only 625 sentences from the much smaller Hartmann et al.
+To do so, we require suitable training data from both document types, and a robust evaluation setup.
+Training Data
+As our FrAC corpus is too small to support fine-tuning, we derive large-scale labeled data sets from the MFC (news articles) and Hartmann et al.
+(2019) (comments).
+We sentence-split both datasets and derive sentence-level annotations from the original span-level annotations: we first map the original frame labels to our 9 classes; then we identify spans with matching labels from at least two annotators with a minimum three-word overlap.
+Following previous methods (Naderi and Hirst, 2017; Hartmann et al., 2019; Park et al., 2022), agreed spans were mapped to the sentence-level: sentences fully covered by agreed spans inherited the span labels.
+This results in approximately 63,000 labeled sentences predominantly from the MFC (including 625 sentences from the much smaller Hartmann et al.
 (2019) dataset).
@@
 Frame Classification Models
-To automatically classify frames in our news articles and comments, we fine-tuned three models — RoBERTa-Large (Liu et al., 2019b) (full fine-tuning), LLaMA3.1-8B-Instruct (Grattafiori et al., 2024), and Qwen2.5-7B-Instruct (Qwen et al., 2025) (both parameter-efficient fine-tuning) — using the sentence-level annotated data described above on a random 80/10/10 (training/dev/test) split, stratified by both frame and topic.
+To automatically classify frames in news articles and comments, we fine-tuned three models – RoBERTa-Large (Liu et al., 2019b) (full fine-tuning), LLaMA3.1-8B-Instruct (Grattafiori et al., 2024), and Qwen2.5-7B-Instruct (Qwen et al., 2025) (both parameter-efficient fine-tuning) – using the sentence-level annotated data described above with a random 80/10/10 (training/dev/test) split, stratified by frame and topic.
 Fine-tuning parameters are reported in Appendix A, and prompting details in Appendix B.
 RoBERTa Llama Qwen
-In-Domain 0.66 0.65 0.64
-Table 2: Macro-F1 scores of three fine-tuned LMs on MFC + Hartmann data.
+P R F1 P R F1 P R F1
+0.70 0.65 0.66 0.68 0.65 0.65 0.64 0.64 0.64
+Table 3: Precision (P), recall (R) and Macro-F1 scores for the three fine-tuned LMs on MFC + Hartmann data.
+Label P R F1
+Economic 0.77 0.83 0.80
+Morality 0.70 0.70 0.70
+Fairness and Equality 0.68 0.59 0.64
+Legality and Crime 0.83 0.86 0.85
+Political and Policy 0.77 0.85 0.81
+Security and Defense 0.69 0.59 0.63
+Health and Safety 0.73 0.80 0.76
+Cultural Identity 0.71 0.54 0.61
+Public Opinion 0.61 0.61 0.61
+Other 0.53 0.14 0.22
+Table 4: Per-frame precision, recall, and Macro-F1 scores for fine-tuned RoBERTa on MFC + Hartmann data.
 X-Domain FrAC (A) FrAC (C)
-RoBERTa 0.59 0.77 0.83
-Table 3: Macro-F1 scores of Lama cross-domain (X-Domain) and on FrAC articles (A) and comments (C).
+P R F1 P R F1 P R F1
+0.65 0.59 0.59 0.74 0.81 0.77 0.84 0.85 0.83
+Table 5: Precision, Recall and Macro-F1 scores of RoBERTa cross-domain (X-Domain) and on FrAC articles (A) and comments (C).
 Evaluation
@@
 Table 6: SOCC and NYT data statistics by topic and source, including the average number of comments per article, after the pre-processing explained in Section 6.
-Table 2 shows little difference in the models’ performance on the 10% MFC/Hartmann test split described above, with RoBERTa-Large slightly outperforming the larger Qwen2.5-Instruct and LLaMA3.1-8B-Instruct.
-This result aligns with recent literature on frame classification, where BERT-based pre-trained models have consistently demonstrated good performance (Sajwani et al., 2024; Daffara et al., 2025).
-Since RoBERTa also requires substantially less fine-tuning time (approximately 20 minutes vs.
-13 hours for PEFT-tuning of LlaMA and Qwen), all remaining experiments in this paper are based on the fine-tuned RoBERTa model.
-To evaluate model generalizability to unseen topics, we train RoBERTa on all but one MFC topics, evaluating on the held-out topic, averaging over all five combinations.
-Table 3 (X-Domain) revealed an expected, albeit relatively small drop in RoBERTa performance (from 0.66 In-Domain in Table 2 to 0.59), suggesting a reasonable ability to generalize to unseen topics, which is important because the model will be applied to topics not present in the MFC training set.
-Finally, we evaluated RoBERTa on the annotated FrAC corpus (Table 3, FrAC), i.e., a data set from different outlets, time points and distinct yet overlapping set of topics (Section 5 and Table 4).
+The averaged results in Table 3 show little performance difference across models, with RoBERTa-Large slightly outperforming the larger Qwen2.5-Instruct and LLaMA3.1-8B-Instruct.
+This result aligns with recent literature on frame classification, where BERT-based pre-trained models have consistently demonstrated strong performance (Sajwani et al., 2024; Daffara et al., 2025).
+Since RoBERTa also requires substantially less fine-tuning time (approximately 20 minutes versus 13 hours for PEFT-tuning of LlaMA and Qwen), all remaining experiments in this paper are based on fine-tuned RoBERTa.
+Table 4 reports the per-label breakdown for the best performing model, RoBERTa.
+To evaluate model generalizability to unseen topics, we train RoBERTa on all but one MFC topics, evaluate on the held-out topic, average over all five combinations.
+Table 5 (X-Domain) revealed an expected, albeit relatively small drop in RoBERTa performance (from 0.66 In-Domain in Table˜4 to 0.59), suggesting a reasonable ability to generalize to unseen topics, which is important because the model will be applied to topics not present in the MFC training set.
+Finally, we evaluated RoBERTa on the annotated FrAC corpus (Table 5, FrAC), i.e., a data set that differs from the MFC fine-tuning data in its outlets, time points and distinct yet overlapping set of topics (Section˜5 and Table˜6).
 The model achieved strong performance on FrAC with an F1 score of 0.77 on articles and 0.83 on comments, demonstrating effective generalization to reader comments despite being trained primarily on formal news articles.
+[^9] FrAC was annotated directly at the sentence level using the consolidated 9+1 frame taxonomy, whereas MFC labels were created post-hoc from span-level annotations, introducing label noise which likely negatively affected model performance.
 This result validates our central methodological contribution: that sentence-level frame classification can successfully bridge the gap between formal news articles and informal reader comments.
-Having validated RoBERTa’s performance on our gold-standard FrAC corpus, we then applied it to the full sentence-tokenized SOCC and NYT data sets (Section 4.1) for our large-scale framing effects analysis.
-[^9] We used the NLTK library https://www.nltk.org/ for sentence tokenization of article and comment texts.
+Having validated RoBERTa’s performance on our gold-standard FrAC corpus, we then applied it to the full sentence-tokenized SOCC and NYT data sets (Section˜4.1) for our large-scale framing effects analysis.
+[^10] We used the NLTK library (https://www.nltk.org/) for sentence tokenization of article and comment texts.
 6.2 Dominant Frame Detection
-We analyze framing effects between articles and comments in terms of their respective dominant frames.
+We analyze framing effects between articles and comments in terms of their respective dominant frames as well as article secondary frames.
 We define the dominant frame as the most frequent sentence-level frame label in a given document.
+Secondary frames are all other frames that were assigned to at least one sentence.
 To reduce noise from near ties, we only assign a dominant frame if the most frequent frame meets both of the following criteria: it appears in at least 3 sentences and it represents at least 40% of all sentences in the document.
@@
 6.3 Topic Extraction
-We applied BERTopic (Grootendorst, 2022) separately to SOCC and NYT to extract topics from news articles, assigning each article a probability distribution across all identified topics.
+We apply BERTopic (Grootendorst, 2022) separately to SOCC and NYT to extract topics from news articles, assigning each article a probability distribution across all identified topics.
 [^11] We report the full list of parameters used in Appendix C.
@@
 The extracted topics showed substantial thematic overlap across data sets, enabling cross-dataset comparison.
-We manually identified eleven overarching, semantically cohesive topics which we use in our downstream analysis (Table 4).
+We manually identified eleven overarching, semantically cohesive topics which we use in our downstream analysis (Table˜6).
 Evaluation
@@
 Two authors of the paper independently reviewed each assigned topic and judged whether it was acceptable or not.
-We observed a high degree of inter-annotator agreement (96%).
+We observe a high degree of inter-annotator agreement (96%).
 The overall accuracy of the model’s predicted topics, calculated against the labels assigned by the annotators was found to be 93% (on both comments and articles).
 6.4 Resulting Data Set
-With our framework and validated implementation at hand, we next apply it to analyze framing effects across the SOCC and NYT data sets (Section 4.1) and eleven selected topics (Section 6.3).
+With our framework and validated implementation at hand, we next apply it to analyze framing effects across the SOCC and NYT data sets (Section˜4.1) and eleven selected topics (Section˜6.3).
 We only include article-comment pairs where both share the same predicted topic (excluding off-topic comments to study framing effects when discussions remain on-topic) and where dominant frames could be extracted for both articles and comments.
-We also filter out comments shorter than 5 words, and repeated texts.
+We also filter out comments shorter than 5 words, and duplicate comments.
 This final data set consists of nearly 3,000 articles and over 380,000 associated comments, as detailed in Table 6.
 7 Framing Effects across Topics and Outlets
-Using our corpus of articles and comment frames from two outlets, annotated with both dominant frames and labels, we next address our three research questions (Section 3) in turn.
-Figure 2: News article frames are retained in reader comments to varying degrees (X- and Y-axes).
-The degree of retention correlates strongly across two media outlets: New York Times (NYT) and The Globe and Mail (SOCC).
-7.1 To What Extent are Frames Retained vs Changed?
-We quantitatively investigated the dominant frames employed in both the news articles and corresponding reader comments across the New York Times (NYT; USA) and The Globe and Mail (SOCC; Canada), analyzing the extent to which the dominant frame of the article was retained in the associated reader comments.
+We now address our three research questions (Section 3) asking to what extent dominant article frames stick with readers?
+How do these trends vary by topic?
+And, if the dominant frame does not stick, how readers reconstruct the media frame in their comments?
+Figure 2: Predicted frame retention rates in reader comments based on regression-derived marginal effects.
+X- and Y-axes show predicted retention for SOCC and NYT, respectively.
+Retention varies by frame type, and the predictions correlate strongly across the two media outlets.
+7.1 To what extent are dominant article frames retained in reader comments?
+In line with our definition of frame retainment as the echoing of the dominant frame of a news article, we first investigate how often the dominant frame kept in its associated direct readers comments.
+We do so across the New York Times (NYT; USA) and The Globe and Mail (SOCC; Canada).
 Overall, we find that NYT readers demonstrate stronger frame retention (50.7%) compared to SOCC readers (37.3%).
-Conversely, at least half of the comments (almost two thirds in SOCC) change away from the dominant article frame indicating that readers often do not retain the original article framing, but change the message.
+Conversely, at least half of the comments (almost two thirds in SOCC) change away from the dominant article frame indicating that readers often do not retain the original primary framing of the article.
 In section 7.3, we analyze these reframing patterns in depth.
+We next investigate how this pattern varies by frame type.
+We do so through a logistic regression mixed effects model that predicts frame retention (i.e., comment frame being identical to the dominant article frame) as a function of fixed effects for dominant article frame and article topic, and a random effect controlling for article ID, [^12] We control for article ID due to repeated measurements, given that we have multiple comments per article.
+$y_{i}\sim\text{topic}_{d}+\text{frame}_{d}+(1|\text{id}_{d}),$ (1)
+where the input is a document-response pair $i=(d,r)$, $y_{i}$ is a binary indicator of frame retention ($\text{frame}_{d}{==}\text{frame}_{r}$).
+Figure 3: Predicted topic retention rates in reader comments based on regression-derived marginal effects.
+Points above the diagonal line indicate higher predicted retention in NYT, while points below indicate higher predicted retention in SOCC.
+Topic names are shortened for readability.
+Here, we are interested in the marginal effect of $\text{frame}_{d}$ on $y$ as derived from the fitted model coefficients.
 Figure 2 displays the frame retention patterns separately for each dominant article frame.
 Looking at the spread of frames along the x- and y-axis independently, we can see that the extent of retention varies by frame.
-Specifically, frames such as "Political and Policies", "Economic", and "Health and Safety" show high retention rates.
-In contrast, "Morality", "Fairness and Equality", and "Cultural Identity" exhibit substantially lower retention rates, suggesting that readers are more likely to reframe content that is presented through frames that require alignment with personal beliefs or values to be accepted (Chong and Druckman, 2007).
-[…]
+Specifically, frames such as “Political and Policies”, “Economic”, and “Health and Safety” show high retention rates.
+In contrast, “Public Opinion”, “Fairness and Equality”, and “Cultural Identity” exhibit substantially lower retention rates.
+This suggests that readers are more likely to reframe content that is presented through frames that require alignment with personal beliefs or values to be accepted (Chong and Druckman, 2007).
+To formally test whether the frame choices in reader comments were influenced by the frames used in the corresponding articles – that is, whether the distribution of frames in comments is not independent of the article’s dominant frame – we conducted chi-squared tests of independence.
+For each source frame $f$, we tested the null hypothesis that the presence of $f$ in an article is independent of the presence of $f$ in a comment.
+For all frames, we were able to reject the null hypothesis of independence ($p\ll.001$), although with small to moderate effect sizes (e.g., small effects: “Public Opinion” Cramér’s V = 0.065 in NYT; moderate effects: “Economic” V = 0.357 in SOCC).
 Figure 2 additionally reveals that the reframing patterns are remarkably consistent across NYT and SOCC, i.e., all points cluster close to the diagonal line of perfect correlation.
-This suggests that our uncovered patterns might reveal a general trend in news reader behavior, rather than being idiosyncratic to a particular outlet or sample, although future analyses across more diverse outlets and countries will need to verify this trend.
-Figure 3: Frame retention rates by topic across news outlets.
-Points above the diagonal line indicate higher retention in NYT, while points below indicate higher retention in SOCC.
-Topic names are shortened for enhanced visibility.
-7.2 Do Framing Effects Vary Across Topics?
-Here, we explore the effect of topics on framing effects.
-We cannot estimate topic effects from the observed data directly, because the effect of topics is confounded by the topic-specific frame distributions.
-To address this, we estimate the effect of article topic on frame retention using logistic regression mixed effects models.
-Given a document-response pair $i=(d,r)$:
-$y_{i}\sim\text{topic}_{d}+(1|\text{frame}_{d})+(1|\text{id}_{d}),$ (1)
-where $y_{i}$ is a binary indicator of frame retention ($\text{frame}_{d}{==}\text{frame}_{r}$).
-Document (article) topic ($\text{topic}_{d}$) is a fixed effect of topic on frame retention.
-And $(1|\text{frame}_{d})$ and $(1|\text{id}_{d})$ are fixed effects, controlling for the impact of the article frame ($\text{frame}_{d}$) and article id $(\text{id}_{d})$, respectively.
-[^11] We control for article ID due to repeated measurements, given that we have multiple comments per article.
-We then obtain the marginal effects of $\text{topic}_{d}$ on $y$ from the fitted model, removing any influence of the random effects.
-The resulting probabilities of frame retention per topic are shown in Figure 3 for NYT (y-axis) and SOCC (x-axis).
-We find that framing effects vary by topic, but the degree of variance is much higher for NYT compared to SOCC suggesting that these patterns are more outlet specific than the framing-based variation discussed above and shown in Figure 2.
-This offers an explanation for the inconsistent findings in previous single-topic studies.
-Some topics exhibit consistent retention patterns across outlets: Coverage of Trump & Elections demonstrates relatively high retention in both data sets (NYT: 31.3%, SOCC: 42.7%), while Healthcare and Education also show moderate to strong retention.
-International conflict topics are consistently more often reframed by the readers: Israel-Palestine conflict and Syria & IS show the lowest retention rates across both sources, ranging between 18-24%.
-Other topics reveal outlet-specific differences: Immigration shows substantially higher frame retention in NYT than SOCC, as well as the coverage of the Russia-Ukraine conflict.
-7.3 How Do Readers Reframe?
+This cross-outlet consistency addresses a key limitation of previous single-outlet studies and suggests that our uncovered patterns might reveal a general trend in news reader behavior, rather than being idiosyncratic to a particular outlet or sample.
+7.2 How does frame retention vary across topics?
+While previous studies have exclusively focused on a single topic and have often produced contradictory results, here we can more rigorously inspect the variation of framing effects across topics under our unified framework.
+We explore the effect of topics on framing effects using the same regression framework as in Section 7.1, but this time focusing on the marginal effects of topic on retention from equation (1).
+This provides predicted frame retention probabilities per topic, controlling for frame type and article-specific effects.
+The resulting predicted probabilities of frame retention per topic are shown in Figure 3 for NYT (y-axis) and SOCC (x-axis).
+We find that predicted framing effects vary considerably by topic, with notable differences between outlets.
+Several topics show consistent retention patterns across both outlets.
+Trump & Elections, for instance, demonstrates high predicted retention in both outlets.
+In contrast, international conflicts like Israel-Palestine and Syria & IS show consistently low retention across both outlets, clustering in the lower portion of the distribution.
+Other topics reveal substantial outlet-specific differences.
+Immigration and Healthcare show notably higher predicted retention in NYT than SOCC, as does coverage of the Russia-Ukraine conflict and LGBT+ Rights, as shown in Figure 3.
+These outlet-specific variations, which have not been examined in previous works, confirm that "different kinds of issues are interpreted by the media and by the public in different ways" (Neuman et al., 1992, pg.
+17), suggesting that topic-based retention and reframing patterns are more outlet-specific than the frame-based patterns in Figure 2.
+This finding is in line with the inconsistent results reported in previous small-scale, single-topic studies in the social sciences (Sections 2 and 3).
+In this paper, we deliberately remain on the descriptive level and refrain from explaining the observed differences, being cogniscent of many confounding factors that may be at play.
+However, our framework paves the way for future causal analyses of framing across different topics and news outlets.
+7.3 If readers reject the dominant article frame, how do they reframe?
+Prior communication studies have used varied approaches to analyze reframing - from binary annotations of the presence of frames (e.g.,Holton et al.
+(2014); Muñiz et al.
+(2015)) to dominant frame identification (e.g., Zhou and Moy (2007); Koteyko et al.
+(2013)).
+However, none systematically examined how readers reframe when they diverge from the dominant article frame – whether they introduce entirely novel frames or selectively emphasize non-dominant frames already present in the source.
+Here, we address this question directly.
 Comments change away from the dominant article frame 50% (NYT) or 63% (SOCC) of the time, so a natural question is how readers reframe articles.
-Figure 4 summarizes the flow of reframing in the NYT across all topics.
-Readers predominantly reframe issues into political terms, regardless of the original article framing, but a particularly large proportion of value-laden article frames (“Fairness and Equality”, “Public Opinion”) are recast as “Political and Policies”.
-[^12] This does not necessarily mean, though, that readers “politicize” all articles – “Political and Policies” frame is also the most reframed one, and its relative proportion in comments is actually lower than in the articles.
-This pattern, albeit to a lesser degree, also emerges in SOCC (Figure 6 in Appendix), where readers also often recast different frames into “Economy”.
-We also investigated topic-specific patterns of reframing for topics that included at least 100 articles.
-For instance, in the case of Immigration (Table 5), SOCC readers both reframe “Cultural Identity” into “Political and Policies” terms, and vice versa, suggesting a tight interconnection of these two perspectives.
-NYT readers, on the other hand, primarily reframe into ‘harder’ perspectives like “Economic” and “Legality and Crime” and interestingly, among these top five reframings (as shown in the table), the “Economic” framing occurs only on the reader side.
-Table 9 in the Appendix shows the analysis for three additional topics, revealing varying trends across topics and outlets.
-Figure 4: Reframing patterns in the New York Times showing the flow from primary article frames (left) to primary comment frames (right) with proportion of frames in articles and comments, respectively.
-We plot standardized frame proportions in articles, to reduce visual clutter.
-Reframing Pattern N
-SOCC Cultural Identity → Political & Policies 134
-Political & Policies → Cultural Identity 131
-Health & Safety → Cultural Identity 120
-Economic → Political & Policies 108
-Health & Safety → Political & Policies 99
-NYT Legality & Crime → Political & Policies 1388
-Political & Policies → Economic 1311
-Political & Policies → Legality & Crime 1139
-Legality & Crime → Economic 1024
-Political & Policies → Cultural Identity 950
-Table 5: Five most common reframings of the topic ‘Immigration’ in SOCC (top) and NYT (bottom) as ‘dominant article frame $\rightarrow$ dominant comment frame’.
+In our analysis, we distinguish between two types of reframing: (a) complete reframing, where the comment frame does not appear in the source article at all, and (b) selective reframing, where comments adopt secondary frames which are present in the source article, but do not dominate the message.
+In SOCC, 11.7% of comment frames exhibit complete reframing, while 62.8% show selective reframing.
+In NYT, these fractions are 5.2% and 49.3%.
+respectively.
+This shows that, rather than completely accepting or rejecting the dominant source framing, audiences predominantly engage in selective interpretation by adopting secondary frames from the article.
+Figure 4 shows reframing patterns for the NYT.
+Focusing first on complete reframing (Figure˜4(a)), Political framing dominates source articles but commentators introduce diverse alternative frames.
+In particular readers reframe in terms of value-laden frames including “Morality” (increasing from 0.8% in articles to 26.4% in Comments) or “Cultural Identity”, as well as “Economic” framing.
+This pattern holds across both NYT and SOCC (as shown in Appendix˜G).
+Selective reframing patterns in NYT are shown in Figure 4(b).
+Here, the frame distributions in the comments align more closely with the dominant article frame distributions, although we again observe an increase in value-laden frames (“Morality”, “Health and Safety”) as well as “Economic” framing.
+Topic-specific Reframing
+We investigate topic-specific reframing patterns for topics that included at least 100 articles.
+For instance, in the case of Immigration (Table 7, left), SOCC readers engage in selective reframing by emphasizing both “Cultural Identity” frames when “Political and Policies” dominate the article, and vice versa, suggesting a tight interconnection of these two perspectives.
+NYT readers show different selective reframing patterns, primarily shifting from “Legality and Crime” frames to “Political and Policies” frames, and from “Political and Policies” to “Economic frames”, indicating a tendency toward more policy-oriented and economic interpretations of immigration coverage.
+Complete reframing differs substantially (Table 7, right).
+In NYT, complete reframing shows “Economic” frames being introduced when they were entirely absent from articles.
+Again, both outlets reframe more objective (economic, political, legal) views into value-laden frames (cultural, morality, fairness).
+Tables 11 and 12 in the Appendix show selective and complete reframing analysis for additional topics, revealing substantial topic-specific differences.
+In conclusion, the results discussed here align with the constructionist view of framing as "interactive, vulnerable and prone to counter-frames" (Gorp, 2007, pg.
+76), with audiences engaging in topic-specific patterns of selective retention for frames that resonate with them, whilst actively reframing those that do not.
+(a) Complete reframing.
+(b) Selective reframing.
+Figure 4: Reframing in the New York Times.
+(a) Complete reframing patterns where readers use frames that never appear anywhere in the original articles.
+(b) Selective reframing where readers emphasize secondary frames that differ from the dominant article frame but exist elsewhere in the article.
+Left percentages show dominant article frame distributions; right percentages show comment frame distributions among complete (a) / selective (b) reframing cases.
+Selective Reframing Complete Reframing
+SOCC Cultural $\rightarrow$ Political Political $\rightarrow$ Cultural
+Health $\rightarrow$ Cultural Economic $\rightarrow$ Political
+Political $\rightarrow$ Cultural Health $\rightarrow$ Morality
+Health $\rightarrow$ Political Economic $\rightarrow$ Cultural
+Economic $\rightarrow$ Political Political $\rightarrow$ Fairness
+NYT Legality $\rightarrow$ Political Political $\rightarrow$ Economic
+Legality $\rightarrow$ Economic Political $\rightarrow$ Legality
+Political $\rightarrow$ Economic Legality $\rightarrow$ Cultural
+Political $\rightarrow$ Legality Political $\rightarrow$ Cultural
+Political $\rightarrow$ Cultural Political $\rightarrow$ Morality
+Table 7: Reframing “Immigration” in SOCC (top) and NYT (bottom).
+Each cell shows the five most common selective (left) and complete (right) shifts from dominant article $\rightarrow$ comment frame.
+Frame names are abbreviated to their first word.
 8 Discussion
@@
 While existing computational methods treat framing as isolated classification tasks focused solely on source content, we examined the dynamic interaction between media framing and audience reception.
-In uncovering systematicities (RQ1; Figure 4) and topic-specific differences (RQs 2-3) in a large-scale study, we bridged disconnect between large-scale yet isolated NLP approaches and small-scale communication studies that lack generalizability across topics and contexts.
-In applying one framing conceptualization across eleven topics, our findings suggest that inconsistent results in prior work stem from topic- and outlet-specific variation in framing effects.
-We also found consistent trends, such as value-laden frames (“Fairness and Equality”, “Morality”) being more likely to be reframed.
-This is in line with theoretical formulations of framing, according to which individuals more readily accept frames that match their pre-existing values (Chong and Druckman, 2007).
-Our analysis reveals outlet-specific trends — e.g., NYT readers’ political reframing and SOCC readers’ economic focus — likely reflect interactions between editorial stance, audience demographics, and national context.
-While frame-based patterns are relatively stable across outlets, topic effects vary more substantially.
-Since our comparison conflates country (US vs.
-Canada), editorial leaning (center-left vs.
-center-right), and time period (2012–2016 vs.
-2017–2018), we refrain from strong interpretation.
-Future work should disentangle these factors to clarify framing dynamics and inform broader interdisciplinary research.
-A key methodological contribution of our framework lies in demonstrating reliable frame classification across fundamentally different text genres – from formal news articles to informal reader comments.
+In uncovering systematicities (RQ1; Figure˜2) and topic-specific differences (RQs 2-3) in a large-scale study, we bridged the disconnect between large-scale yet isolated NLP approaches and small-scale communication studies that lack generalizability.
+In applying one unified framing conceptualization across eleven topics, our findings suggest that inconsistent results on reframing patterns in prior work can be partially explained by topic- and outlet-specific variation in framing effects.
+We also identified consistent trends, such as value-laden frames (“Fairness and Equality”, “Morality”) being more likely to be reframed (Figure 2), but also more likely to be introduced in the context of complete reframing (Figure 4(a)).
+This divergent pattern suggests that audiences reject value-laden framings that conflict with their worldview while actively introducing those that align with it.
+This extends theoretical formulations about audience personal formulation of messages (Chong and Druckman, 2007) by showing that readers not only selectively accept frames matching their values, but also actively impose their preferred value frames onto discussions where they were originally absent.
+In distinguishing selective from complete reframing, we show that readers predominantly reframe in selective ways, providing empirical evidence for what framing theory has long posited: that news consumers reformulate and engage with content selectively based on their preferences (Neuman et al., 1992; Scheufele, 1999) – a phenomenon that small-scale social science studies could not disentangle systematically, and that NLP studies have overlooked by disregarding response analysis.
+Our analysis reveals that frame-based patterns are relatively stable across outlets, while topic-specific effects vary more widely.
+These differences likely arise from complex interactions between unobservable factors including reader demographics and national context.
+Furthermore, our data sets conflate country (US versus Canada), editorial leaning (center-left NYT versus center-right SOCC), and time period (2012–2016 NYT versus 2017–2018 SOCC).
+We hence deliberately maintain a descriptive level and leave disentanglement of these factors for future work.
+A key methodological contribution lies in demonstrating reliable frame classification across fundamentally different text genres – formal news articles and informal reader comments.
 This cross-genre generalization addresses a critical gap in computational framing analysis, where models typically operate within homogeneous text types.
-Our validation shows that sentence-level frame classification can successfully detect frames in structurally and stylistically different content, with strong performance on our human-annotated data set FrAC.
+Our validation shows strong performance on our human-annotated data set FrAC.
 Our study builds on the MFC, a widely-used framework in NLP.
-Yet, it is by no means the only or the optimal framework, and it has been criticized for the broad nature and lack of theoretical grounding of its frames (Ali and Hassan, 2022).
-We note that our method is framework-agnostic: alternative framing theories could be readily integrated, whether more generic frameworks (e.g., Semetko and Valkenburg (2000)), or issue-specific taxonomies tailored to particular domains, a rich direction for future work.
-Similarly, our methods for dominant frame detection and sentence-level annotation, while supporting discovery of major trends, limit finer-grained framing analyses, which could be explored in future work.
-Finally, our framework holds potential relevance for other NLP tasks including argumentation interpretation and generation, and even disinformation detection.
+Yet, it is by no means the only and optimal frameworks, and it has been criticized for the broad nature and lack of theoricatical ground of its framings (Ali and Hassan, 2022).
+We note that our method is framework-agnostic: alternative framing theories could be readily integrated, whether more generic frameworks (e.g., Semetko and Valkenburg (2000)), or issue-specific taxonomies tailored to particular domains.
+Similarly, our approach supports both large-scale analyses of framing trends and more fine-grained investigations that future work could further explore the role of minority frames.
+Finally, our framework holds potential relevance for other NLP tasks including argumentation interpretation and generation, and disinformation detection.
 The systematic patterns we identify in how audiences reconstruct frames could inform models of persuasive text generation, helping systems understand how different framings might be received and transformed by target audiences.
 9 Conclusion
-We introduced a scalable computational framework for analyzing media framing effects as an interactive process between source content and audience responses, grounded in a constructionist view on framing.
-Through analysis of over 3,000 news articles and 380,000 comments across two outlets and eleven topics, we demonstrate that readers reframe content at least as frequently as they retain original article frames.
-Frame retention varies systematically by frame type, with value-laden frames being more likely to be reframed, and when reframing occurs, readers consistently shift toward political interpretation regardless of the original framing.
-To support our analysis, we contributed a robust framing scheme from an established taxonomy, together with robust classifiers that work across different document types, providing a foundation for future research on media framing effects.
+This work addresses the gap between framing theory and its computational operationalization.
+While framing theory posits that individuals respond differentially to framed information – accepting, rejecting, or modifying frames based on their cognitive schemas – computational approaches have largely reduced framing to taxonomic classification, neglecting audience reception processes.
+We introduced a scalable computational framework for analyzing media framing effects as an interactive process between source content and audience responses.
+Through analysis of over 3,000 news articles and 380,000 comments across two outlets and eleven topics, we demonstrate that readers reframe content at least as frequently as they retain original frames, often reframing selectively by picking up on secondary article frames.
+Frame retention varies systematically by frame type with objective frames (e.g., “Health & Safety”) more likely to be retained than value-laden frames (e.g., “Morality”).
+We contributed a robust framing scheme with classifiers that work across document types, providing a foundation for future research.
+Our analysis leaves open questions for future work, including how framing effects propagate through nested comment conversations and identifying causal linguistic properties that explain specific framing effects, or thread and temporal analysis.
 Acknowledgment
@@
   • Data splitting:
+    • nosep
       Training set: 80%
+    • nosep
       Validation set: 10%
+    • nosep
       Test set: 10%
@@
 Framing Categories
-<Displays list of framing categories as in Table 8>
+See Table 10
 Frame Description Example
@@
 Other If the sentence is not framed in any of the proposed categories, select this label.
-Table 8: Framing Categories used in this work.
+Table 10: Framing categories shown on the annotation interface.
 Practice Phase
@@
 Happy annotating!
-Appendix F Reframing Patterns in SOCC
-Figure 6: Reframing patterns in the SFU Opinion Corpus (SOCC) showing the flow from primary article frames (left) to primary comment frames (right).
-We print the original % of frames in articles and comments, respectively.
-This is reflected on the bar height on the right, but not on the left to reduce visual clutter.
-Appendix G Reframing Patterns by Topic
-Topic SOCC NYT
-Reframing Pattern Count Reframing Pattern Count
-Climate Change Political and Policies → Economic 1126 Political and Policies → Economic 1590
-Economic → Political and Policies 907 Economic → Political and Policies 857
-Political and Policies → Cultural Identity 185 Political and Policies → Morality 690
-Political and Policies → Public Opinion 181 Political and Policies → Health and Safety 686
-Economic → Cultural Identity 156 Political and Policies → Cultural Identity 465
-Immigration Cultural Identity → Political and Policies 134 Legality and Crime → Political and Policies 1388
-Political and Policies → Cultural Identity 131 Political and Policies → Economic 1311
-Health and Safety → Cultural Identity 120 Political and Policies → Legality and Crime 1139
-Economic → Political and Policies 108 Legality and Crime → Economic 1024
-Health and Safety → Political and Policies 99 Political and Policies → Cultural Identity 950
-Syria & IS Security and Defense → Political and Policies 405 Security and Defense → Political and Policies 321
-Political and Policies → Security and Defense 360 Security and Defense → Morality 165
-Security and Defense → Morality 127 Health and Safety → Political and Policies 149
-Security and Defense → Cultural Identity 120 Health and Safety → Morality 123
-Political and Policies → Morality 96 Security and Defense → Cultural Identity 106
-Trump & Elections Political and Policies → Cultural Identity 407 Political and Policies → Morality 2516
-Political and Policies → Public Opinion 293 Political and Policies → Cultural Identity 2373
-Political and Policies → Economic 206 Political and Policies → Health and Safety 1362
-Political and Policies → Fairness and Equality 178 Cultural Identity → Political and Policies 1145
-Cultural Identity → Political and Policies 145 Political and Policies → Economic 1039
-Table 9: Top 5 reframing patterns by topic for SOCC and NYT outlets, showing the transition from one frame to another and the frequency count.
+Appendix F Reframing Patterns by Topic
+Figure 11 and Figure 12 show the 5 most frequent selective and complete reframing patterns in both news outlets in coverage of climate change, immigration, Syria & IS, Trump & elections.
+Topic Outlet Selective Reframing Pattern Count
+$C$ SOCC Political $\to$ Economic 1001
+Economic $\to$ Political 854
+Political $\to$ Public Opinion 113
+Economic $\to$ Cultural 90
+Political $\to$ Cultural 79
+NYT Political $\to$ Economic 1585
+Economic $\to$ Political 857
+Political $\to$ Health 428
+Political $\to$ Cultural 345
+Political $\to$ Legality 232
+IM SOCC Cultural $\to$ Political 132
+Health $\to$ Cultural 113
+Political $\to$ Cultural 99
+Health $\to$ Political 96
+Economic $\to$ Political 85
+NYT Legality $\to$ Political 1388
+Legality $\to$ Economic 955
+Political $\to$ Economic 927
+Political $\to$ Legality 903
+Political $\to$ Cultural 769
+SY SOCC Political $\to$ Security 352
+Security $\to$ Political 326
+Security $\to$ Morality 78
+Security $\to$ Cultural 64
+Political $\to$ Morality 41
+NYT Security $\to$ Political 321
+Health $\to$ Political 146
+Health $\to$ Morality 123
+Security $\to$ Morality 117
+Security $\to$ Cultural 79
+TR SOCC Political $\to$ Cultural 319
+Political $\to$ Public Opinion 256
+Political $\to$ Economic 146
+Cultural $\to$ Political 145
+Political $\to$ Fairness 108
+NYT Political $\to$ Cultural 2136
+Political $\to$ Morality 2070
+Cultural $\to$ Political 1145
+Political $\to$ Health 983
+Political $\to$ Economic 774
+Table 11: Top 5 selective reframing patterns by topic and outlet. $C$ = Climate Change, IM = Immigration, SY = Syria, TR = Trump.
+Topic Outlet Complete Reframing Pattern Count
+$C$ SOCC Political $\to$ Economic 125
+Political $\to$ Cultural 106
+Political $\to$ Health 72
+Political $\to$ Public Opinion 68
+Economic $\to$ Cultural 66
+NYT Political $\to$ Morality 473
+Political $\to$ Health 258
+Political $\to$ Cultural 120
+Health $\to$ Morality 62
+Economic $\to$ Morality 51
+IM SOCC Political $\to$ Cultural 32
+Economic $\to$ Political 23
+Health $\to$ Morality 22
+Economic $\to$ Cultural 17
+Political $\to$ Fairness 17
+NYT Political $\to$ Economic 384
+Political $\to$ Legality 236
+Legality $\to$ Cultural 189
+Political $\to$ Cultural 181
+Political $\to$ Morality 177
+SY SOCC Security $\to$ Political 79
+Security $\to$ Cultural 56
+Political $\to$ Morality 55
+Security $\to$ Morality 49
+Political $\to$ Cultural 39
+NYT Security $\to$ Morality 48
+Security $\to$ Cultural 27
+Security $\to$ Economic 25
+Fairness $\to$ Morality 14
+Health $\to$ Cultural 14
+TR SOCC Political $\to$ Cultural 88
+Political $\to$ Fairness 70
+Political $\to$ Economic 60
+Political $\to$ Legality 60
+Political $\to$ Morality 38
+NYT Political $\to$ Morality 446
+Political $\to$ Health 379
+Political $\to$ Economic 265
+Political $\to$ Legality 252
+Political $\to$ Cultural 237
+Table 12: Top 5 complete reframing patterns by topic and outlet. $C$ = Climate Change, IM = Immigration, SY = Syria, TR = Trump.
+Appendix G Reframing in SOCC
+Figure 6 and Figure 7 show complete and selective reframing patterns in SOCC, for comparison against NYT results.
+Figure 6: Complete reframing in SOCC showing cases where readers use frames never appearing in articles.
+Figure 7: Selective reframing in SOCC showing cases where readers emphasize secondary frames differing from dominant article frames but existing elsewhere in articles.